### v1

# Preconditioning FEM discretisations of the high-frequency Maxwell equations by either perturbing the coefficients or adding absorption

Abstract
We prove bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$ where $\mathsf{A}_{\ell}$, $\ell=1,2$, are the Galerkin matrices corresponding to finite-element discretisations of the time-harmonic Maxwell equations $k^{-2}{\rm curl}(\mu_{\ell}^{-1}{\rm curl}E_{\ell})-\epsilon_{\ell}E_{\ell}=f$; i.e., we consider the situation where the Maxwell FEM matrix is preconditioned by the FEM matrix arising from the same Maxwell problem but with different coefficients. An important special case is when the perturbation consists of adding absorption (in the spirit of “shifted Laplacian preconditioning” for the Helmholtz equation). The results of this paper are the Maxwell analogues of the Helmholtz results in [16, 20], and confirm a conjecture in the recent preprint [26]. These results are obtained by putting the Maxwell problem in an abstract framework that also includes the Helmholtz problem; as a byproduct we weaken the assumptions required to obtain the Helmholtz results in [16, 20].
1 Context and motivation
The design of good preconditioners for the linear systems arising from finite-element discretisations of high-frequency time-harmonic wave problems – such as the Helmholtz or time-harmonic Maxwell equations – is a longstanding open problem; see, e.g., the reviews [9, 12, 18, 17] and the references therein. This paper considers the theory of the following two, related, questions involving preconditioning these linear systems.
Preconditioning using absorption.
The idea of preconditioning discretisations of the Helmholtz equation with (approximations of) discretisations of the same Helmholtz problem with added absorption was introduced in [11, 10]. Seeking to rigorously understand this method, the paper [16] proved bounds on ${\mathsf{I}}-\mathsf{A}_{\varepsilon}^{-1}\mathsf{A}$, where $\mathsf{A}$ is the Helmholtz Galerkin matrix, and $\mathsf{A}_{\varepsilon}$ is the Helmholtz Galerkin matrix with $k^{2}\mapsto k^{2}+{\rm i}\varepsilon$, where $k$ is the (real and large) wavenumber. [16] proved that, for a particular nontrapping Helmholtz problem, $\|{\mathsf{I}}-\mathsf{A}_{\varepsilon}^{-1}\mathsf{A}\|_{2}\leq C\varepsilon/k$. An important question is then whether the action of $\mathsf{A}_{\varepsilon}^{-1}$ can be (provably) efficiently approximated when $\varepsilon\sim k$. This question was investigated for one-level domain-decomposition methods in [21], and the recent work [24] proved that certain hybrid two-level Schwarz preconditioners (with problem-adapted basis functions in the coarse space) formed with absorption $\varepsilon\sim k$ give a $k$-independent number of GMRES iterations when applied to the original Helmholtz problem.
“Nearby preconditioning”.
The results of [16] were generalised in [20] to bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$, where $\mathsf{A}_{\ell}$, $\ell=1,2$, are the Galerkin matrices of $k^{-2}\nabla\cdot(A_{\ell}\nabla u_{\ell})+n_{\ell}u_{\ell}=f$. The motivation for studying this situation comes from uncertainty quantification (UQ): calculating quantities of interest for the solution of the Helmholtz equation $k^{-2}\nabla\cdot(A\nabla u)+nu=f$ with random coefficients requires the solution of many deterministic Helmholtz problems, each with different coefficients $A$ and $n$. Bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$ then indicate to what extent a previously-calculated inverse of one of the Galerkin matrices can be used as a preconditioner for other Galerkin matrices; see [32, §4.6] and the recent work [35] for UQ algorithms using this idea.
Content and motivation for the present paper.
The present paper
  • proves the Helmholtz results of [16, 20] under less-restrictive assumptions than in [16, 20], and
  • generalises these results to the time-harmonic Maxwell equations.
Our main motivation is the recent work [26]: this paper gives the first frequency-explicit analysis of a two-level domain-decomposition method for the high-frequency time-harmonic Maxwell equations, under the assumption that the Helmholtz results of [16, 20] hold also for the time-harmonic Maxwell equations; the present paper shows that this assumption indeed holds.
Outline of the paper.
§2 states the abstract assumptions under which the main results are proved, and §3 describes Helmholtz and Maxwell problems that are covered by these abstract assumptions. §4 recaps properties of the inf-sup constant of a sesquilinear form and its $k$-dependence for Helmholtz and Maxwell. §5 states the main results, and §6 proves them.
2 Abstract assumptions
${\cal H}\subset{\cal H}_{0}$ are Hilbert spaces with $\|\cdot\|_{{\cal H}_{0}}\leq\|\cdot\|_{{\cal H}}$. We identify ${\cal H}_{0}$ with its dual so that ${\cal H}\subset{\cal H}_{0}\subset{\cal H}^{*}$. The linear operator ${\mathcal{D}}:{\cal H}\to{\cal H}_{0}$ with $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$. $b(\cdot,\cdot)$ is a continuous sesquilinear form on ${\cal H}$,
$a_{\ell}(u,v):=\big{(}\mu_{\ell}^{-1}{\mathcal{D}}u,{\mathcal{D}}v\big{)}_{{\cal H}_{0}}+b(u,v)-\big{(}\epsilon_{\ell}u,v\big{)}_{{\cal H}_{0}},\quad\ell=1,2,$ (2.1)
where $\mu_{\ell}^{-1}:{\cal H}_{0}\to{\cal H}_{0}$ and $\epsilon_{\ell}:{\cal H}_{0}\to{\cal H}_{0}$ are bounded linear operators for $\ell=1,2$. Finally for $\ell=1,2,$ there exists $C_{\rm G1,\ell},C_{\rm G2,\ell}>0$ such that
$\big{|}a_{\ell}(v,v)+C_{\rm G2,\ell}\left\|v\right\|^{2}_{{\cal H}_{0}}\big{|}\geq C_{\rm G1,\ell}\left\|v\right\|^{2}_{{\cal H}}\quad\text{ for all }v\in{\cal H}.$ (2.2)
We make four remarks.
  • In all examples below, $\|v\|^{2}_{{\cal H}}$ equals $\|{\mathcal{D}}v\|_{{\cal H}_{0}}^{2}+\|v\|_{{\cal H}_{0}}^{2}$ plus possibly an additional term; thus the condition $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$ is immediately satisfied.
  • These assumptions imply that $a_{\ell}(\cdot,\cdot)$ is continuous; i.e., there exist $C_{\rm cont,\ell}>0$ such that
    $\big{|}a_{\ell}(u,v)\big{|}\leq C_{\rm cont,\ell}\left\|u\right\|_{{\cal H}}\left\|v\right\|_{{\cal H}}\quad\text{ for all }u,v\in{\cal H}.$
  • If $a(\cdot,\cdot)$ satisfies the Gårding inequality
    $\Re a_{\ell}(v,v)\geq C_{\rm G1,\ell}\left\|v\right\|^{2}_{{\cal H}}-C_{\rm G2,\ell}\left\|v\right\|^{2}_{{\cal H}_{0}}\quad\text{ for all }v\in{\cal H},$ (2.3)
    then (2.2) holds (since $|z|\geq\Re z$ for all $z\in\mathbb{C}$).
  • If $a_{\ell}(\cdot,\cdot)$ satisfies these assumptions, then so does
    $a_{\ell}^{*}(u,v):=\overline{a_{\ell}(v,u)}=\big{(}\mu_{\ell}^{*}{\mathcal{D}}u,{\mathcal{D}}v\big{)}_{{\cal H}_{0}}+\overline{b(v,u)}-\big{(}\epsilon_{\ell}^{*}u,v\big{)}_{{\cal H}_{0}},$
    with the same $C_{\rm G1,\ell},C_{\rm G2,\ell}$ (this is because $|a_{\ell}(v,v)+C_{\rm G2,\ell}\left\|v\right\|^{2}_{{\cal H}_{0}}|=\big{|}\overline{a_{\ell}(v,v)}+C_{\rm G2,\ell}\left\|v\right\|^{2}_{{\cal H}_{0}}\big{|}$).
3 Helmholtz and Maxwell problems satisfying the abstract assumptions
3.1 Helmholtz problems
In all the Helmholtz problems, $\Omega$ is a bounded Lipschitz open set, ${\cal H}_{0}=L^{2}(\Omega)$ (scalar valued), ${\mathcal{D}}:=k^{-1}\nabla$, and
$\left\|v\right\|^{2}_{{\cal H}}:=\left\|k^{-1}\nabla v\right\|^{2}_{L^{2}(\Omega)}+\left\|v\right\|^{2}_{L^{2}(\Omega)}$
(so that immediately $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$). In addition $\mu_{\ell}^{-1}$ are bounded, symmetric, matrix functions on $\Omega$ with $\Re\mu_{\ell}^{-1}>0$ (in the sense of quadratic forms) and $\epsilon_{\ell}$ are bounded, scalar-valued functions on $\Omega$. Note that these properties of $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ are satisfied by the coefficients coming from a radial PML by, e.g., [14, Lemma 2.3].
3.1.1 Helmholtz problem with Dirichlet boundary conditions on part of the boundary (including PML approximation to scattering by a Dirichlet or Neumann obstacle)
:=${\cal H}:=\big{\{}u\in H^{1}(\Omega)\,:\,\text{ $u=0$ on at least one connected component of $\partial\Omega$}\big{\}}$
and $b(\cdot,\cdot)=0$. In this case, the sesquilinear form corresponds to the Helmholtz operator
$-k^{-2}\nabla\cdot(\mu_{\ell}^{-1}\nabla)-\epsilon_{\ell},$ (3.1)
with zero Dirichlet boundary conditions on at least one connected component of $\partial\Omega$ and zero Neumann boundary conditions on the rest.
3.1.2 Helmholtz problem with impedance boundary conditions on part of the boundary
With $\Gamma_{\rm imp}$ a non-empty connected component of $\partial\Omega$, and $\Gamma_{\rm D}$ a (possibly empty) connected component of $\partial\Omega$ that is not equal to $\Gamma_{\rm imp}$,
:=${\cal H}:=\big{\{}v\in H^{1}(\Omega)\,:\,v=0\text{ on }\Gamma_{\rm D}\big{\}},$
and
$b(u,v):=-{\rm i}k^{-1}\big{(}\theta u,v\big{)}_{L^{2}(\Gamma_{\rm imp})}$
for some real-valued $\theta\in L^{\infty}(\Gamma_{\rm imp})$ such that ${\rm essinf}_{\Gamma_{\rm imp}}\theta>0$. To see that (2.2) is satisfied, observe that $a(\cdot,\cdot)$ satisfies the Gårding inequality (2.3) (with the term on $\Gamma_{\rm imp}$ playing no role).
In this case, the sesquilinear form corresponds to the Helmholtz operator (3.1) with the impedance boundary condition
$k^{-1}\partial_{n,\mu^{-1}}u-{\rm i}\theta u=g_{\rm imp}\quad\text{ on }\Gamma_{\rm imp}$
$u=0$ on $\Gamma_{\rm D}$, and $k^{-1}\partial_{n,\mu^{-1}}u=g_{\rm N}$ on $\partial\Omega\setminus(\Gamma_{\rm D}\cup\Gamma_{\rm imp})$ (where the data $g_{\rm imp}$ and $g_{\rm N}$ depend on the given right-hand side). Here $\partial_{n,\mu^{-1}}$ is the conormal derivative; recall that this is such that $\partial_{n,\mu^{-1}}u:=n\cdot\mu^{-1}\nabla u$ when $u\in H^{2}(\Omega)$, where $n$ is the outward-pointing unit normal vector to $\partial\Omega$ (see, e.g., [28, Lemmas 4.2 and 4.3]).
3.1.3 Helmholtz problem with exact Dirichlet-to-Neumann map imposed on the boundary of a ball
$\Omega=\{x:|x|\leq R\}\setminus\overline{\Omega_{-}}$, where $\Omega_{-}$ is a bounded Lipschitz open set with connected open complement (so that the scattering problem with obstacle $\Omega_{-}$ makes sense). ${\cal H}$ equals either $H^{1}(\Omega)$ or $\{v\in H^{1}(\Omega)\,:\,v=0\text{ on }\partial\Omega_{-}\}$.
$b(u,v):=-{\rm i}k^{-1}\big{(}{\rm DtN}_{k}u,v\big{)}_{L^{2}(\Gamma_{\rm imp})},$
where ${\rm DtN}_{k}$ is the exact Dirichlet-to-Neumann map for the Helmholtz equation in the exterior of $\{x:|x|\leq R\}$; see, e.g., [1, Equations 3.5 and 3.6] for explicit expressions in terms of Fourier series (in 2-d) or spherical harmonics (in 3-d) and Bessel and Hankel functions. To see that (2.2) is satisfied, recall that $\Re(-{\rm i}{\rm DtN}_{k})\geq 0$ by, e.g., [1, Corollary 3.1], so that $a(\cdot,\cdot)$ satisfies the Gårding inequality (2.3).
3.2 Maxwell problems
In all the Maxwell problems $\Omega$ is a bounded Lipschitz open set, ${\cal H}_{0}=L^{2}(\Omega)$ (vector valued), ${\mathcal{D}}:=k^{-1}{\rm curl}$, and the sesquilinear form corresponds to the Maxwell operator
$k^{-2}{\rm curl}\big{(}\mu^{-1}_{\ell}{\rm curl}\big{)}-\epsilon_{\ell}.$ (3.2)
3.2.1 Maxwell problem with PEC boundary condition (including PML approximation to scattering by a PEC obstacle)
${\cal H}:=H_{0}({\rm curl};\Omega)=\big{\{}v\in H({\rm curl};\Omega):v_{T}=0\big{\}}$
where $v_{T}$ is the tangential trace such that $v_{T}:=(n\times v)\times n$ for smooth $v$, where $n$ is the outward-pointing unit normal vector (see, e.g., [31, Equation 3.46 and Theorem 3.31]).
$\left\|v\right\|^{2}_{{\cal H}}:=\left\|k^{-1}{\rm curl}\,v\right\|^{2}_{L^{2}(\Omega)}+\left\|v\right\|^{2}_{L^{2}(\Omega)}$
and $b(\cdot,\cdot)=0$. $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ are bounded, symmetric, matrix functions on $\Omega$ with $\Re\mu_{\ell}^{-1}>0$ and $\Re\epsilon_{\ell}>0$ (in the sense of quadratic forms). (Note that these properties are satisfied by the coefficients coming from a radial PML by, e.g., [3, Appendix A].)
3.2.2 Maxwell problem with impedance boundary conditions on part of the boundary
With $\Gamma_{\rm imp}$ a non-empty connected component of $\partial\Omega$,
:=${\cal H}:=\big{\{}u\in H^{1}(\Omega)\,:\,\text{ $u_{T}=0$ on $\partial\Omega\setminus\Gamma_{\rm imp}$ and $u_{T}\in L^{2}(\Gamma_{\rm imp})$}\big{\}},$
$\left\|v\right\|^{2}_{{\cal H}}:=k^{-2}\left\|{\rm curl}\,v\right\|^{2}_{L^{2}(\Omega)}+\left\|v\right\|^{2}_{L^{2}(\Omega)}+k^{-1}\left\|v_{T}\right\|^{2}_{L^{2}(\Gamma_{\rm imp})},$
and
$b(u,v):=-{\rm i}k^{-1}\big{(}\theta u_{T},v_{T}\big{)}_{L^{2}(\Gamma_{\rm imp})},$
for some real-valued $\theta\in L^{\infty}(\Gamma_{\rm imp})$ such that ${\rm essinf}_{\Gamma_{\rm imp}}\theta>0$. $\mu_{\ell}^{-1}$ are real-valued bounded, symmetric, matrix functions on $\Omega$ with $\mu_{\ell}^{-1}>0$ (in the sense of quadratic forms). $\epsilon_{\ell}$ are real-valued, bounded, symmetric matrix functions on $\Omega$.
To see that (2.2) is satisfied, observe that, since $\mu^{-1}_{\ell}$ and $\epsilon_{\ell}$ are all real, $\Im(a(v,v))$ consists only of $\Im(b(v,v))$, which is bounded below by $k^{-1}({\rm essinf}_{\Gamma_{\rm imp}}\theta)\|v\|^{2}_{L^{2}(\Gamma_{\rm imp})}$.
In this case, the sesquilinear form corresponds to the Maxwell operator (3.2) with the impedance boundary condition
$k^{-1}\big{(}\mu_{\ell}^{-1}{\rm curl}\,u\big{)}\times n-{\rm i}\theta u_{T}=g\quad\text{ on }\Gamma_{\rm imp}$
(where $g$ depends on the given right-hand side) and the PEC boundary condition $u_{T}=0$ on $\Gamma\setminus\Gamma_{\rm imp}$.
Remark 3.1 (The reason for the particular weighting with $k$).
Most papers on the numerical analysis of the Helmholtz and time-harmonic Maxwell equations write these equations as $(\Delta+k^{2})u=f$ and $({\rm curl}\,{\rm curl}-k^{2})E=f$ (in the constant coefficient case) and use the norms $\|v\|^{2}_{H^{1}}=\|\nabla v\|^{2}_{L^{2}}+k^{2}\|v\|^{2}_{L^{2}}$ and $\|w\|_{H({\rm curl})}^{2}=\|{\rm curl}w\|^{2}_{L^{2}}+k^{2}\|w\|^{2}_{L^{2}}$. The advantage of rescaling the PDEs and norms so that every derivative appears with a $k^{-1}$ is that none of the constants in §2 (i.e. $C_{\rm G1,\ell},C_{\rm G2,\ell},$ and $C_{\rm cont,\ell}$) depend on $k$, and (consequentially) the $k$-dependence of the norm of the solution operator is the same between any two spaces for which the solution operator is defined; i.e., the ${\cal H}_{0}\to{\cal H}_{0}$ norm, the ${\cal H}_{0}\to{\cal H}$ norm, and the ${\cal H}^{*}\to{\cal H}$ norm all have the same $k$-dependence – see Lemma 4.2 below.
4 Recap of properties of the inf-sup constant of $a_{\ell}(\cdot,\cdot)$
Let ${\mathcal{A}}_{\ell}:{\cal H}\to{\cal H}^{*}$ be the operator associated to $a_{\ell}(\cdot,\cdot)$; i.e., $\langle{\mathcal{A}}_{\ell}u,v\rangle_{{\cal H}^{*}\times{\cal H}}=a_{\ell}(u,v)$ for all $u,v\in{\cal H}$. We recall the following standard result.
Theorem 4.1 (Inf-sup condition equivalent to bound on inverse operator).
The conditions
$\inf_{u\in{\cal H}\setminus\{0\}}\sup_{v\in{\cal H}\setminus\{0\}}\frac{\big{|}a_{\ell}(u,v)\big{|}}{\left\|u\right\|_{{\cal H}}\left\|v\right\|_{{\cal H}}}\geq\gamma^{-1}\text{ and for all }v\in{\cal H}\setminus\{0\},\sup_{u\in{\cal H}\setminus\{0\}}|a_{\ell}(u,v)|>0$ (4.1)
and
$\big{\|}{\mathcal{A}}_{\ell}^{-1}\big{\|}_{{\cal H}^{*}\to{\cal H}}\leq\gamma$
are equivalent.
References for the proof.
See, e.g., [22, Lemma 6.5.3], [33, Theorem 2.1.44]. The first bound in (4.1) is that $\|{\mathcal{A}}_{\ell}u\|_{{\cal H}^{*}}\geq\gamma^{-1}\|u\|_{{\cal H}}$ for all $u\in{\cal H}$, and the second bound in (4.1) implies that the kernel of ${\mathcal{A}}_{\ell}^{*}$ is empty, and thus the image of ${\mathcal{A}}_{\ell}$ equals ${\cal H}^{*}$; see, e.g., [28, Theorem 2.13(iii)]. ∎
The following lemma, proved in §6.4 below using the Gårding-type inequality (2.2), shows that all norms of ${\mathcal{A}}_{\ell}^{-1}$ are equivalent, with constants depending only on $C_{\rm G1,\ell}$ and $C_{\rm G2,\ell}$ (and hence independent of $k$).
Lemma 4.2 ($k$-independent equivalence of norms of ${\mathcal{A}}_{\ell}^{-1}$).
If ${\mathcal{A}}_{\ell}^{-1}:{\cal H}^{*}\to{\cal H}$ exists, then
$\big{\|}\mathcal{A}_{\ell}^{-1}\big{\|}_{{\cal H}_{0}\to{\cal H}}\leq\big{\|}\mathcal{A}_{\ell}^{-1}\big{\|}_{{\cal H}^{*}\to{\cal H}}\leq(C_{\rm G1,\ell})^{-1}\Big{(}1+C_{\rm G2,\ell}\big{\|}\mathcal{A}_{\ell}^{-1}\big{\|}_{{\cal H}_{0}\to{\cal H}}\Big{)}$ (4.2)
and
$\big{\|}\mathcal{A}_{\ell}^{-1}\big{\|}_{{\cal H}_{0}\to{\cal H}_{0}}\leq\big{\|}\mathcal{A}_{\ell}^{-1}\big{\|}_{{\cal H}_{0}\to{\cal H}}\leq(C_{\rm G1,\ell})^{-1/2}\big{\|}\mathcal{A}_{\ell}^{-1}\big{\|}_{{\cal H}_{0}\to{\cal H}_{0}}\sqrt{C_{\rm G2,\ell}+\big{\|}\mathcal{A}_{\ell}^{-1}\big{\|}_{{\cal H}_{0}\to{\cal H}_{0}}^{-1}}.$ (4.3)
Remark 4.3 (The $k$-dependence of $\|{\mathcal{A}}^{-1}_{\ell}\|$ for Helmholtz and Maxwell).
When $\mu_{\ell}$ and $\epsilon_{\ell}$ are both constant multiples of the identity in part of the domain, $\|{\mathcal{A}}^{-1}_{\ell}\|\geq Ck$ – this can be proved by considering data that is a cut-off function multiplied by a plane wave; see, e.g., [1, Lemma 3.10] (for Helmholtz) and [4, §1.4.1], [30, Example 3.4] (for Maxwell).
For the scattering problem, if the scatterer is nontrapping then $\|{\mathcal{A}}^{-1}_{\ell}\|\leq Ck$; this has been proved in a wide variety of Helmholtz cases (see the literature review in [19]) and for the Maxwell problem with certain nontrapping coefficients in [4] or a nontrapping PEC obstacle in [36, §2]. Under trapping, $\|{\mathcal{A}}^{-1}_{\ell}\|\gg k$; see the results and literature reviews in [25], [2].
For the PML problem, [13, Theorem 1.6] proved that the norm of the solution operator of the Helmholtz radial PML problem is bounded by the norm of the solution operator of the corresponding Helmholtz scattering problem; we expect that the same result holds for the Maxwell PML problem – indeed, this result was recently proved for the case of no scatterer and Cartesian PML in [6, Lemma 10].
5 Statement of the main results
5.1 Notation for the Galerkin matrices
Let ${\cal H}_{N}\subset{\cal H}$ be a finite-dimensional space with basis $\{\phi_{j}\}_{j=1}^{N}$. Let
$\mathsf{A}_{\ell}=\mathsf{S}_{\mu^{-1}_{\ell}}+\mathsf{B}-\mathsf{M}_{\epsilon_{\ell}},\quad\ell=1,2,$ (5.1)
where
$(\mathsf{S}_{\mu^{-1}_{\ell}})_{ij}=\big{(}\mu^{-1}_{\ell}{\mathcal{D}}\phi_{j},{\mathcal{D}}\phi_{i}\big{)}_{{\cal H}_{0}},\quad(\mathsf{M}_{\epsilon_{\ell}})_{ij}=\big{(}\epsilon_{\ell}\phi_{j},\phi_{i}\big{)}_{{\cal H}_{0}},\quad\text{ and }\quad\mathsf{B}_{ij}=b(\phi_{j},\phi_{i}).$ (5.2)
Let the matrix $\mathsf{D}$ be such that
$\left\|v_{N}\right\|_{{\cal H}}^{2}=\left\|\mathbf{V}\right\|_{\mathsf{D}}^{2}\quad\text{ for }v_{N}=\sum_{j=1}^{N}V_{j}\phi_{j}\,\text{ with }\,V_{j}\in\mathbb{C}^{N}.$ (5.3)
We also use the notation $\|\cdot\|_{\mathsf{D}}$ to denote the induced norm on matrices. Finally, let $\|\cdot\|_{2}$ denote the Euclidean norm on $\mathbb{C}^{N}$, and let $m_{\pm}$ be such that
$m_{-}\|\mathbf{V}\|_{2}\leq\|v_{N}\|_{{\cal H}_{0}}\leq m_{+}\|\mathbf{V}\|_{2}\quad\text{ for all }v_{N}\in{\cal H}_{N}.$ (5.4)
Remark 5.1.
With $\mathsf{M}$ the mass matrix, i.e., $(\mathsf{M})_{ij}=(\phi_{j},\phi_{i})_{{\cal H}_{0}}$, (5.4) is equivalent to the bounds
$m_{-}^{2}\|\mathbf{V}\|_{2}^{2}\leq\big{(}\mathsf{M}\mathbf{V},\mathbf{V}\big{)}_{2}\leq m_{+}^{2}\|\mathbf{V}\|_{2}^{2}\quad\text{ for all }\mathbf{V}\in\mathbb{C}^{N};$
i.e., the quantity $(m_{+}/m_{-})^{2}$ (whose square root appears in the bound (5.8) below) is the ratio of the largest to the smallest eigenvalue – i.e., the condition number – of the mass matrix $\mathsf{M}$.
5.2 The main result: bounds on $\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\|$ and $\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\|$
Theorem 5.2 (Bounds on $\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\|$ and $\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\|$).
Suppose that the assumptions of §2 hold. Suppose that $({\cal H}_{N})_{N=1}^{\infty}$ are such that there exists $C_{1}>0$ such that, for all $N\in\mathbb{Z}^{+}$,
$\inf_{u_{N}\in{\cal H}_{N}\setminus\{0\}}\sup_{v_{N}\in{\cal H}_{N}\setminus\{0\}}\frac{\big{|}a_{1}(u_{N},v_{N})\big{|}}{\left\|u_{N}\right\|_{{\cal H}}\left\|v_{H}\right\|_{{\cal H}}}\geq\frac{1}{C_{1}\big{\|}{\mathcal{A}}_{1}^{-1}\big{\|}_{{\cal H}^{*}\to{\cal H}}}$ (5.5a)
and
[…]
(i.e., the discrete inf-sup constant of $a_{1}(\cdot,\cdot)$ is bounded below by a constant times the continuous inf-sup constant (4.1)). If
$\Big{(}\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big{)}C_{1}\big{\|}{\mathcal{A}}^{-1}_{1}\big{\|}_{{\cal H}^{*}\to{\cal H}}\leq 1/2$ (5.6)
then $\mathsf{A}_{2}^{-1}$ exists and
$\max\Big{\{}\big{\|}{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big{\|}_{\mathsf{D}},\big{\|}{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big{\|}_{\mathsf{D}^{-1}}\Big{\}}$
$\hskip 85.35826pt\leq 2\Big{(}\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big{)}C_{1}\big{\|}{\mathcal{A}}^{-1}_{1}\big{\|}_{{\cal H}^{*}\to{\cal H}}.$ (5.7)
Furthermore, if $\mu_{1}=\mu_{2}$, then
$\max\Big{\{}\big{\|}{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big{\|}_{2},\big{\|}{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big{\|}_{2}\Big{\}}\leq 2\frac{m_{+}}{m_{-}}\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{1}\big{\|}{\mathcal{A}}^{-1}_{1}\big{\|}_{{\cal H}^{*}\to{\cal H}}.$ (5.8)
We make the following four immediate remarks.
  • (i)
    The assumptions of Theorem 5.2 – by design – involve $a_{1}(\cdot,\cdot)$ and not $a_{2}(\cdot,\cdot)$; i.e., the assumptions are on the problem that we want to solve (involving $\mathsf{A}_{1}$), rather than the problem used for preconditioning (involving $\mathsf{A}_{2}$).
  • (ii)
    Since ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}=\mathsf{A}_{2}^{-1}(\mathsf{A}_{2}-\mathsf{A}_{1})$, the right-hand sides of both (5.7) and (5.8) should be understood as $\|\mathsf{A}_{2}-\mathsf{A}_{1}\|\|\mathsf{A}_{2}^{-1}\|$; i.e., the norm of the perturbation times the norm of the solution operator. The key point is that the conditions (5.5a) and (5.6) mean that $\|\mathsf{A}_{2}^{-1}\|$ is bounded by $\|{\mathcal{A}}_{2}^{-1}\|_{{\cal H}^{*}\to{\cal H}}$, which in turn is bounded by $\|{\mathcal{A}}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}$.
  • (iii)
    §5.3 gives sufficient conditions for (5.5a) to hold (including a new result on this); i.e., conditions under which the discrete inf-sup constant is bounded by the continuous inf-sup constant.
  • (iv)
    In all our Helmholtz and Maxwell examples, the ${\cal H}_{0}\to{\cal H}_{0}$ norms of the coefficient differences appearing on the right-hand sides of both (5.7) and (5.8) are bounded by their $L^{\infty}$ norms (since $\|\mu v\|_{L^{2}(\Omega)}\leq\|\mu\|_{L^{\infty}(\Omega)}\left\|v\right\|_{L^{2}(\Omega)}$ for all $v\in L^{2}(\Omega)$).
Corollary 5.3 (Theorem 5.2 specialised to preconditioning with absorption).
Suppose that the assumptions of §2 hold with $\mu_{2}=\mu_{1}$ and $\epsilon_{2}=(1+{\rm i}\alpha)\epsilon_{1}$ with $\alpha>0$. Suppose that $({\cal H}_{N})_{N=1}^{\infty}$ are such that there exists $C_{1}>0$ such that (5.5) hold for all $N\in\mathbb{Z}^{+}$. If
$\alpha C_{1}\big{\|}{\mathcal{A}}^{-1}_{1}\big{\|}_{{\cal H}^{*}\to{\cal H}}\leq 1/2$
then $\mathsf{A}_{2}^{-1}$ exists,
$\max\Big{\{}\big{\|}{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big{\|}_{\mathsf{D}},\big{\|}{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big{\|}_{\mathsf{D}^{-1}}\Big{\}}\leq 2\alpha C_{1}\big{\|}{\mathcal{A}}^{-1}_{1}\big{\|}_{{\cal H}^{*}\to{\cal H}}$ (5.9)
and
$\max\Big{\{}\big{\|}{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big{\|}_{2},\big{\|}{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big{\|}_{2}\Big{\}}\leq 2\frac{m_{+}}{m_{-}}\alpha C_{1}\big{\|}{\mathcal{A}}^{-1}_{1}\big{\|}_{{\cal H}^{*}\to{\cal H}}.$ (5.10)
Remark 5.4 (Recovering the Helmholtz results of [16] about preconditioning with absorption).
The set up in [16] considers preconditioning the Helmholtz equation with absorption by letting $k^{2}\mapsto k^{2}+{\rm i}\varepsilon$. In Corollary 5.3 this corresponds to setting $\alpha:=\varepsilon/k^{2}$. By Remark 4.3, $\big{\|}{\mathcal{A}}^{-1}_{1}\big{\|}_{{\cal H}^{*}\to{\cal H}}\sim k$ when the Helmholtz problem is nontrapping. The right hand side of (5.10) is then proportional to $\varepsilon/k$, which is the result proved in [16, Theorem 1.4].
Remark 5.5 (Implications of Theorem 5.2 for preconditioning).
If the right-hand side of (5.7) is $\leq c<1$, then
  • the preconditioned fixed point iteration $\mathbf{x}^{n+1}=\mathbf{x}^{n}+\mathsf{A}_{2}^{-1}({\bf b}-\mathsf{A}_{1}\mathbf{x}^{n})$ for solving $\mathsf{A}_{1}\mathbf{x}={\bf b}$ satisfies
    $\left\|\mathbf{x}-\mathbf{x}^{n}\right\|_{\mathsf{D}}\leq c^{n}\left\|\mathbf{x}-\mathbf{x}^{0}\right\|_{\mathsf{D}},$ (5.11)
  • when GMRES is applied in the $\mathsf{D}$ inner product, the residual $\mathbf{r}^{n}:=\mathsf{A}_{2}^{-1}{\bf b}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\mathbf{x}^{n}$ satisfies
    $\frac{\left\|\mathbf{r}^{n}\right\|_{\mathsf{D}}}{\left\|\mathbf{r}^{0}\right\|_{\mathsf{D}}}\leq\left(\frac{2\sqrt{c}}{(1+c)^{2}}\right)^{n}$ (5.12)
    (as a consequence of the Elman estimate [8, 7]; see [16, Corollary 1.9]/[20, Corollary 5.5]).
Similar statements hold for right-preconditioning and/or the fixed-point iteration/GMRES in the Euclidean inner product (using (5.8) instead of (5.7)).
We highlight that, at least in the case of preconditioning by absorption, the action of $\mathsf{A}_{2}^{-1}$ must be further approximated to provide a practical preconditioner (as discussed in §1).
Remark 5.6 (The analogue of Theorem 5.2 at the level of sesquilinear forms).
A byproduct of the proof of Theorem 5.2 below is that if
$\Big{(}\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big{)}\big{\|}{\mathcal{A}}^{-1}_{1}\big{\|}_{{\cal H}^{*}\to{\cal H}}\leq 1/2$ (5.13)
then ${\mathcal{A}}_{2}^{-1}$ exists and
$\max\Big{\{}\big{\|}I-{\mathcal{A}}_{2}^{-1}{\mathcal{A}}_{1}\big{\|}_{{\cal H}},\big{\|}I-{\mathcal{A}}_{1}{\mathcal{A}}_{2}^{-1}\big{\|}_{{\cal H}}\Big{\}}$
$\hskip 85.35826pt\leq 2\Big{(}\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big{)}\big{\|}{\mathcal{A}}^{-1}_{1}\big{\|}_{{\cal H}^{*}\to{\cal H}}$ (5.14)
(see Remark 6.3). For the Helmholtz case with $\mu_{1}=\mu_{2}$, [20, Lemma 2.7] exhibits $\epsilon_{1},\epsilon_{2}$ such that the bound (5.14) is sharp in its $k$-dependence (in this example $\|{\mathcal{A}}^{-1}_{1}\|_{{\cal H}^{*}\to{\cal H}}\sim k$).
Remark 5.7 (Measuring the difference in the coefficients in weaker norms).
By using standard finite-element inverse estimates, [20, Lemma 5.3] proved (for the Helmholtz case) bounds on $\big{\|}{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big{\|}_{\mathsf{D}}$, $\big{\|}{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big{\|}_{\mathsf{D}}$, $\big{\|}{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big{\|}_{2}$, and $\big{\|}{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big{\|}_{2}$ with $L^{q}$ norms of both $\mu^{-1}_{1}-\mu^{-1}_{2}$ and $\epsilon_{1}-\epsilon_{2}$ appearing on the right-hand sides. These bounds also hold for the Helmholtz and Maxwell problems in §3; however, we do not include them here since (i) these bounds involves powers of $h^{-1}$ (from the inverse estimates) and are thus worse than (5.7) and (5.8), and (ii) the bounds (5.7) and (5.8) are sufficient for studying preconditioning with absorption (becoming (5.9) and (5.10) in Corollary 5.3) and, in particular, verifying the assumption about this in [26].
Remark 5.8 (Theorem 5.2 improves the Helmholtz results of [16] and [20] in three ways).
(i) Theorem 5.2 covers all the Helmholtz problems in §3.1, whereas [16] and [20] only consider the exterior Dirichlet problem with a nontrapping obstacle (with the obstacle allowed to be empty). For this problem, [16] considers truncation of the exterior domain by an impedance boundary condition, and [20] considers truncation by either an impedance boundary condition or the exact Dirichlet-to-Neumann map.
(ii) Theorem 5.2 makes assumptions only about $a_{1}(\cdot,\cdot)$ (i.e., the problem we want to precondition), whereas [16] and [20] make assumptions about both $a_{1}(\cdot,\cdot)$ and $a_{2}(\cdot,\cdot)$; e.g., the mesh threshold in [16, Equation 1.12] is for the problem with absorption, i.e., $a_{2}(\cdot,\cdot)$, and [20, Theorems 2.2 and 2.3] make assumptions on both $a_{1}(\cdot,\cdot)$ and $a_{2}(\cdot,\cdot)$.
(iii) Theorem 5.2 is proved under the discrete inf-sup condition (5.5) (which, by Theorem 4.1, is equivalent to the Galerkin matrix being invertible), whereas the bounds of [16] and [20] are proved under stronger assumptions than a discrete inf-sup condition. Indeed, [16, Equations 1.13 and 1.14] prove bounds analogous to (5.8) under the assumption that, given any data and a sequence of finite-dimensional subspaces, the sequence of Galerkin solutions is quasi-optimal, with constant independent of $k$ – we see in Corollary 5.10 below that this implies that the Galerkin matrix is invertible, with the norm of its inverse bounded by the norm of the PDE solution operator. Furthermore, [20, Lemma 5.3] proves bounds analogous to (5.7) and (5.8) under assumptions weaker than quasi-optimality, but stronger than invertibility of the Galerkin matrix; see [20, Condition 4.2].
5.3 Conditions under which the discrete inf-sup constant is proportional to the (continuous) inf-sup constant
The main observation in this subsection (which we couldn’t find elsewhere in the literature) is that the argument bounding $\|\mathcal{A}_{\ell}^{-1}\|_{{\cal H}^{*}\to{\cal H}}$ in terms of $\|\mathcal{A}_{\ell}^{-1}\|_{{\cal H}_{0}\to{\cal H}}$ (in Lemma 4.2) also holds at the discrete level, giving the following result.
Lemma 5.9 (Bound on discrete inf-sup constant via bound on Galerkin solution).
Suppose that the assumptions in §2 hold and there exists $C_{\rm sol,\ell}>0$ such that, for all $f\in{\cal H}_{0}$, the solution $u_{N}\in{\cal H}_{N}$ to
$a_{\ell}(u_{N},v_{N})=(f,v_{N})_{{\cal H}_{0}}\quad\text{ for all }v_{N}\in{\cal H}_{N}$
exists and satisfies
$\left\|u_{N}\right\|_{{\cal H}}\leq C_{\rm sol,\ell}\left\|f\right\|_{{\cal H}_{0}}.$ (5.15)
Then
$\inf_{u_{h}\in{\cal H}_{h}\setminus\{0\}}\sup_{v_{h}\in{\cal H}_{h}\setminus\{0\}}\frac{\big{|}a_{\ell}(u_{h},v_{h})\big{|}}{\left\|u_{h}\right\|_{{\cal H}}\left\|v_{h}\right\|_{{\cal H}}}\geq\frac{1}{(C_{\rm G1,\ell})^{-1}\big{(}1+C_{\rm G2,\ell}C_{\rm sol,\ell}\big{)}}.$
Several analyses of the $h$-FEM for Helmholtz and Maxwell prove the bound (5.15) directly; e.g., this is done in [27, Theorem 4.1] for the Maxwell impedance problem in §3.2.2 when $\Omega$ a $C^{2}$ domain, $\mu_{\ell}=\epsilon_{\ell}=I$, $({\cal H}_{N})_{N=0}^{\infty}$ consists of Nédélec edge elements of the second type, and $(kh)^{2}k$ sufficiently small.
We now give two other ways of proving the bound (5.15), and hence bounding the discrete inf-sup constant. The first of these is in the asymptotic regime; i.e., when the sequence of Galerkin solutions is quasi-optimal – for the $h$-FEM applied to Helmholtz/Maxwell, this is when $(kh)^{p}\|{\mathcal{A}}_{\ell}^{-1}\|$ is sufficiently small; the second of these is in the preasymptotic regime.
Corollary 5.10 (Bound on discrete inf-sup constant via quasi-optimality).
Suppose that the assumptions in §2 hold and there exists $0<C_{\rm qo,\ell}<\infty$ such that the following holds. Suppose that $({\cal H}_{N})_{N=0}^{\infty}$ is a sequence of finite-dimensional subspaces such that, for all $N$, given $f\in{\cal H}_{0}$ the Galerkin solution $u_{N}$ to
$a_{\ell}(u_{N},v_{N})=(f,v_{N})_{{\cal H}_{0}}\quad\text{ for all }v_{N}\in{\cal H}_{N}$
exists, is unique, and satisfies
$\left\|u-u_{N}\right\|_{{\cal H}}\leq C_{\rm qo,\ell}\left\|(I-\Pi_{N})u\right\|_{{\cal H}},$
where $\Pi_{N}:{\cal H}\to{\cal H}_{N}$ is the orthogonal projection (in the ${\cal H}$ norm) and $u$ is the solution to $a_{\ell}(u,v)=(f,v)_{{\cal H}_{0}}$ for all $v\in{\cal H}$.
Then
$\inf_{u_{h}\in{\cal H}_{h}\setminus\{0\}}\sup_{v_{h}\in{\cal H}_{h}\setminus\{0\}}\frac{\big{|}a_{\ell}(u_{h},v_{h})\big{|}}{\left\|u_{h}\right\|_{{\cal H}}\left\|v_{h}\right\|_{{\cal H}}}\geq\frac{1}{(C_{\rm G1,\ell})^{-1}\Big{(}1+C_{\rm G2,\ell}\big{(}1+C_{\rm qo,\ell}\big{)}\big{\|}\mathcal{A}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\Big{)}}.$
Corollary 5.11 (Bound on discrete inf-sup constant in preasymptotic regime).
Suppose that given $k_{0}>0$ there exists $c$ such that $\big{\|}{\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\geq c$ for all $k\geq k_{0}$. Given $p\in\mathbb{Z}^{+}$, let $({\cal H}_{h})_{h>0}$ be a sequence of finite-dimensional subspaces of ${\cal H}$.
Suppose there exist $C_{j},j=1,2,3,$ such that if
$(kh)^{2p}\big{\|}{\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\leq C_{1}$ (5.16)
then, for all $u\in{\cal H},u_{h}\in{\cal H}_{h}$ satisfying $a_{\ell}(u-u_{h},v_{h})=0$ for all $v_{h}\in{\cal H}_{h}$,
$\left\|u-u_{h}\right\|_{{\cal H}}\leq C_{2}\Big{(}1+(kh)^{p}\big{\|}{\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\Big{)}\left\|(I-\Pi_{h})u\right\|_{\cal H},$ (5.17)
where $\Pi_{h}:{\cal H}\to{\cal H}_{h}$ is the orthogonal projection (in the ${\cal H}$ norm), and, furthermore
$\big{\|}(I-\Pi_{h}){\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\leq C_{3}\Big{(}kh+(kh)^{p}\big{\|}{\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\Big{)}.$ (5.18)
Then, given $0<\varepsilon\leq C_{1}$ there exists $C_{4}>0$ such that if $(kh)^{2p}\big{\|}{\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\leq\varepsilon$ then
$\inf_{u_{h}\in{\cal H}_{h}\setminus\{0\}}\sup_{v_{h}\in{\cal H}_{h}\setminus\{0\}}\frac{\big{|}a_{\ell}(u_{h},v_{h})\big{|}}{\left\|u_{h}\right\|_{{\cal H}}\left\|v_{h}\right\|_{{\cal H}}}\geq\frac{1}{(C_{\rm G1,\ell})^{-1}\Big{(}1+C_{\rm G2,\ell}(1+C_{2}C_{3}C_{4})\big{\|}\mathcal{A}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\Big{)}}.$ (5.19)
Since the proofs of Corollaries 5.10 and 5.11 are so short, we give them here.
Proof of Corollary 5.10.
By the triangle inequality and the fact that $\|(I-\Pi_{N})u\|_{{\cal H}}\leq\left\|u\right\|_{\cal H}$ (since $\Pi_{N}:{\cal H}\to{\cal H}_{N}$ is the orthogonal projection),
$\left\|u_{N}\right\|_{{\cal H}}\leq\left\|u\right\|_{{\cal H}}+\left\|u-u_{N}\right\|_{{\cal H}}\leq(1+C_{\rm qo,\ell})\left\|u\right\|_{{\cal H}}\leq(1+C_{\rm qo,\ell})\left\|{\mathcal{A}}^{-1}_{\ell}\right\|_{{\cal H}_{0}\to{\cal H}}\left\|f\right\|_{{\cal H}_{0}},$
and the result follows from Lemma 5.9. ∎
Proof of Corollary 5.11.
Given $f\in{\cal H}_{0}$, let $u\in{\cal H}$ be the solution to $a_{\ell}(u,v)=(f,v)_{{\cal H}_{0}}$ for all $v\in{\cal H}$. Now, by the combination of (5.17) and (5.18),
$\left\|u-u_{h}\right\|_{{\cal H}}$ $\leq C_{2}\Big{(}1+(kh)^{p}\big{\|}{\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\Big{)}C_{3}\Big{(}kh+(kh)^{p}\big{\|}{\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\Big{)}\left\|f\right\|_{{\cal H}_{0}}.$
Since $(kh)^{2p}\big{\|}{\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\leq\varepsilon$ and $\big{\|}{\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\geq c$, there exists $C_{4}>0$ (depending on $c$ and $\varepsilon$) such that
$\left\|u-u_{h}\right\|_{{\cal H}}\leq C_{2}C_{3}C_{4}\big{\|}{\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\left\|f\right\|_{{\cal H}_{0}}.$
By the triangle inequality, (5.15) holds with
$C_{\rm sol,\ell}:=(1+C_{2}C_{3}C_{4})\big{\|}{\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}},$
and the result (5.19) follows from Lemma 5.9. ∎
Remark 5.12 (${\mathcal{A}}^{-1}_{\ell}$ vs $({\mathcal{A}}^{*}_{\ell})^{-1}$ in the bound (5.18)).
The bound (5.18) is usually proved with ${\mathcal{A}}^{-1}_{\ell}$ replaced by $({\mathcal{A}}^{*}_{\ell})^{-1}$ (see, e.g., [29, 5, 15]). However, if
$a(u,v)=a(\overline{v},\overline{u})\quad\text{ for all }u,v\in{\cal H},$ (5.20)
then the definitions of ${\mathcal{A}}^{*}_{\ell}$ and ${\mathcal{A}}_{\ell}$ imply that $({\mathcal{A}}^{*}_{\ell})^{-1}f=\overline{{\mathcal{A}}_{\ell}^{-1}\overline{f}}$, and thus the bound (5.18) is equivalent to the corresponding bound with ${\mathcal{A}}^{-1}_{\ell}$ replaced by $({\mathcal{A}}^{*}_{\ell})^{-1}$. The condition (5.20) holds for all the Helmholtz and Maxwell problems in §3 either when $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ are Hermitian or when $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ corresponding to a radial PML (see [14, Lemma 2.3]).
Furthermore, the proof of [15, Theorem 1.7], which establishes (5.18) with ${\mathcal{A}}^{-1}_{\ell}$ replaced by $({\mathcal{A}}^{*}_{\ell})^{-1}$ for general Helmholtz problems, also proves (5.18).
Remark 5.13 (Link to other results in the literature).
The result that the discrete inf-sup constant is bounded by the continuous inf-sup constant when $\|(I-\Pi_{h})({\mathcal{A}}^{*}_{\ell})^{-1}\|_{{\cal H}_{0}\to{\cal H}}$ is small is [29, Theorem 4.2]. By the Schatz argument [34], [29, Theorem 4.3], if $\|(I-\Pi_{h})({\mathcal{A}}^{*}_{\ell})^{-1}\|_{{\cal H}_{0}\to{\cal H}}$ is small, then the sequence of Galerkin solutions is quasi-optimal, and so [29, Theorem 4.2] is morally equivalent to Corollary 5.10.
Bounding the discrete inf-sup constant for the $h$-FEM applied to each of the Helmholtz and Maxwell problems in §3.
Suppose that the coefficients and domain satisfy the natural regularity requirements for using degree $p$ polynomials; i.e., the domain is $C^{p,1}$ and the coefficients are piecewise $C^{p-1,1}$.
  • All the Helmholtz problems in §3 satisfy the assumptions of Corollary 5.11 by [15, Theorems 1.5 and 1.7] (with [15, Theorem 1.5] giving (5.17) and [15, Theorem 1.7] giving (5.18)).
  • For the Maxwell PEC problem of §3.2.1 discretised using Nédélec edge elements of either first or second type, the bound (5.17) under the condition (5.16) is proved in [3, Theorem 1.3]. The bound (5.18) is not proved in [3] (the arguments of [3] use a duality argument that “builds in” the splitting used to prove (5.18)). Therefore, although we expect that the bound (5.15) holds in the preasymptotic regime for the Maxwell PEC problem (via a result analogous to Corollary 5.11), right now [3, Theorem 1.3] only allows us to use Corollary 5.10 in the asymptotic regime (i.e. under the condition that $(kh)^{p}\|{\mathcal{A}}^{-1}_{\ell}\|$ is sufficiently small).
  • As noted above, for the Maxwell impedance problem of §3.2.2, the bound (5.15) is proved directly in [27, Theorem 4.1] for the Maxwell impedance problem in §3 when $\Omega$ a $C^{2}$ domain, $\mu_{\ell}=\epsilon_{\ell}=I$, $({\cal H}_{N})_{N=0}^{\infty}$ consists of Nédélec edge elements of the second type, and $(kh)^{2}k$ sufficiently small (note that $\|{\mathcal{A}}^{-1}_{\ell}\|\sim k$ here by [23]).
6 Proofs of the main results
6.1 Bounds on $\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\|$ and $\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\|$ in terms of the discrete inf-sup constant
The heart of the proof of Theorem 5.2 is the following result.
Lemma 6.1.
Suppose that the assumptions of §2 hold. Given $N>0$, if exists $0<C_{{\rm dis},N,2}<\infty$ such that
$\inf_{u_{N}\in{\cal H}_{N}\setminus\{0\}}\sup_{v_{N}\in{\cal H}_{N}\setminus\{0\}}\frac{\big{|}a_{2}(u_{N},v_{N})\big{|}}{\left\|u_{N}\right\|_{{\cal H}}\left\|v_{H}\right\|_{{\cal H}}}\geq\frac{1}{C_{{\rm dis},N,2}}$ (6.1a)
and
[…]
then $\mathsf{A}_{2}^{-1}$ exists and
$\max\Big{\{}\big{\|}{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big{\|}_{\mathsf{D}},\big{\|}{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big{\|}_{\mathsf{D}^{-1}}\Big{\}}\leq\Big{(}\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big{)}C_{{\rm dis},N,2}.$ (6.2)
Furthermore, if $\mu_{1}=\mu_{2}$, then
$\max\Big{\{}\big{\|}{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big{\|}_{2},\big{\|}{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big{\|}_{2}\Big{\}}\leq\frac{m_{+}}{m_{-}}\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}.$ (6.3)
Given Lemma 6.1, to prove Theorem 5.2 we only need to show that the condition (5.5) implies that (6.1) holds with $C_{{\rm dis},N,2}$ bounded above by $\|{\mathcal{A}}^{-1}_{1}\|_{{\cal H}^{*}\to{\cal H}}$.
Proof of Lemma 6.1.
We first show how the bounds on ${\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}$ follow from those on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$. Given a matrix $\mathsf{C}$, let $\mathsf{C}^{\dagger}$ be the conjugate transpose of $\mathsf{C}$ (i.e. the adjoint with respect to $(\cdot,\cdot)_{2}$). Then
$\big{\|}{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big{\|}_{2}=\big{\|}{\mathsf{I}}-(\mathsf{A}_{2}^{\dagger})^{-1}(\mathsf{A}_{1}^{\dagger})\big{\|}_{2}.$
Furthermore, by direct calculation,
$\frac{(\mathbf{V}_{1},\mathsf{C}\mathbf{V}_{2})_{\mathsf{D}^{-1}}}{\left\|\mathbf{V}_{1}\right\|_{\mathsf{D}^{-1}}\left\|\mathbf{V}_{2}\right\|_{\mathsf{D}^{-1}}}=\frac{(\mathsf{C}^{\dagger}\mathbf{W}_{1},\mathbf{W}_{2})_{\mathsf{D}}}{\left\|\mathbf{W}_{1}\right\|_{\mathsf{D}}\left\|\mathbf{W}_{2}\right\|_{\mathsf{D}}}\quad\text{ for all }\mathbf{V}_{j}\in\mathbb{C}^{N},$
where $\mathbf{W}_{j}:=\mathsf{D}^{-1}\mathbf{V}_{j}$, $j=1,2,$. Therefore
$\big{\|}{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big{\|}_{\mathsf{D}^{-1}}=\big{\|}{\mathsf{I}}-(\mathsf{A}_{2}^{\dagger})^{-1}(\mathsf{A}_{1}^{\dagger})\big{\|}_{\mathsf{D}}.$
Recall from §2 that if $a_{\ell}(\cdot,\cdot)$ satisfies the assumptions in §2, then so does $a^{*}_{\ell}(\cdot,\cdot)$ defined by $a^{*}_{\ell}(u,v)=\overline{a_{\ell}(v,u)}$, with the same constants $C_{\rm G1,\ell}$ and $C_{\rm G2,\ell}$. Furthermore, if the discrete inf-sup condition (5.5) on $a(\cdot,\cdot)$ holds, then it also holds for $a^{*}(\cdot,\cdot)$, with the same $C_{{\rm dis},N,2}$; see, e.g., [22, Lemma 6.5.3], [33, Remark 2.1.45] (this is just the result that the norm of the adjoint operator equals the norm of the original operator).
Therefore, once the bound on $\big{\|}{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big{\|}_{\mathsf{D}}$ in (5.7) is established, this bound also holds $\mathsf{A}_{2}$ replaced by $\mathsf{A}_{2}^{\dagger}$ and $\mathsf{A}_{1}$ replaced by $\mathsf{A}_{1}^{\dagger}$, so that
$\big{\|}{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big{\|}_{\mathsf{D}^{-1}}$ $=\big{\|}{\mathsf{I}}-(\mathsf{A}_{2}^{\dagger})^{-1}(\mathsf{A}_{1}^{\dagger})\big{\|}_{\mathsf{D}}$
$\leq\Big{(}\left\|(\mu^{*}_{1})^{-1}-(\mu^{*}_{2})^{-1}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon^{*}_{1}-\epsilon^{*}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big{)}C_{{\rm dis},N,2}$
$=\Big{(}\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big{)}C_{{\rm dis},N,2}.$
Identical reasoning obtains the bound on $\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\|_{2}$ from that on $\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\|_{2}$.
We now prove the bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$ in (5.7) and (5.8). By Theorem 4.1, the discrete inf-sup condition (5.5) implies that $\mathsf{A}_{2}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ is invertible (see also, e.g., [33, Theorem 4.2.1]). Then, by the definitions of $\mathsf{A}_{\ell}$ (5.1), $\mathsf{S}_{\mu}$, and $\mathsf{M}_{\epsilon}$ (5.2),
${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}=\mathsf{A}_{2}^{-1}(\mathsf{A}_{2}-\mathsf{A}_{1})=\mathsf{A}_{2}^{-1}\Big{(}\mathsf{S}_{\mu^{-1}_{2}}-\mathsf{S}_{\mu^{-1}_{1}}-\mathsf{M}_{\epsilon_{2}}+\mathsf{M}_{\epsilon_{1}}\Big{)}=\mathsf{A}_{2}^{-1}\Big{(}\mathsf{S}_{\mu^{-1}_{2}-\mu^{-1}_{1}}-\mathsf{M}_{\epsilon_{2}-\epsilon_{1}}\Big{)}.$
Therefore, to prove (6.2), it is sufficient to prove that
$\big{\|}\mathsf{A}_{2}^{-1}\mathsf{S}_{\mu^{-1}}\big{\|}_{\mathsf{D}}\leq\left\|\mu^{-1}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}\quad\text{ and }\quad\big{\|}\mathsf{A}_{2}^{-1}\mathsf{M}_{\epsilon}\big{\|}_{\mathsf{D}}\leq\left\|\epsilon\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}.$
and to prove (6.3), it is sufficient to prove that
$\big{\|}\mathsf{A}_{2}^{-1}\mathsf{M}_{\epsilon}\big{\|}_{2}\leq\frac{m_{+}}{m_{-}}\left\|\epsilon\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}.$
It is therefore sufficient to prove that, given $\mathbf{F}\in\mathbb{C}^{N}$, the solutions $\mathbf{U}$ and $\mathbf{W}$ to
$\mathsf{A}_{2}\mathbf{U}=\mathsf{S}_{\mu^{-1}}\mathbf{F}\quad\text{ and }\quad\mathsf{A}_{2}\mathbf{W}=\mathsf{M}_{\epsilon}\mathbf{F}$ (6.4)
satisfy
$\left\|\mathbf{U}\right\|_{\mathsf{D}}\leq\left\|\mu^{-1}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}\left\|\mathbf{F}\right\|_{\mathsf{D}},\quad\left\|\mathbf{W}\right\|_{\mathsf{D}}\leq\left\|\epsilon\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}\left\|\mathbf{F}\right\|_{\mathsf{D}},$ (6.5)
and
$m_{-}\left\|\mathbf{W}\right\|_{2}\leq m_{+}\left\|\epsilon\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}\left\|\mathbf{F}\right\|_{2}.$ (6.6)
Given $\mathbf{F}\in\mathbb{C}^{N}$, let $\widetilde{f}\in{\cal H}$ and $\widetilde{F}\in{\cal H}^{*}$ be defined by
$\widetilde{f}:=\sum_{j=1}^{N}F_{j}\phi_{j}\quad\text{ and }\quad\widetilde{F}(v)=\big{(}\mu^{-1}{\mathcal{D}}\widetilde{f},{\mathcal{D}}v\big{)}_{{\cal H}_{0}}.$ (6.7)
The solutions $u_{N}$ and $w_{N}$ to the variational problems
$a_{2}(u_{N},v_{N})=\widetilde{F}(v_{N})\quad\text{ and }\quad a_{2}(w_{N},v_{N})=(\widetilde{f},v_{N})_{{\cal H}_{0}}\quad\text{ for all }v_{N}\in{\cal H}_{N}$ (6.8)
are then such that
$u_{N}=\sum_{j=1}^{N}U_{j}\phi_{j}\quad\text{ and }\quad w_{N}=\sum_{j=1}^{N}W_{j}\phi_{j},$ (6.9)
with $\mathbf{U}$ and $\mathbf{W}$ the solutions to (6.4).
Now, by the bound (6.1a) involving $C_{{\rm dis},N,2}$, the definition of $u_{N}$ (6.8), the definition of $\widetilde{F}$ (6.7), and the fact that $\left\|{\mathcal{D}}\right\|_{{\cal H}\to{\cal H}_{0}}\leq 1$,
$\left\|u_{N}\right\|_{{\cal H}}\leq C_{{\rm dis},N,2}\sup_{v_{N}\in{\cal H}_{N}}\frac{\big{|}a_{2}(u_{N},v_{N})\big{|}}{\left\|v_{N}\right\|_{{\cal H}}}$ $\leq C_{{\rm dis},N,2}\sup_{v_{N}\in{\cal H}_{N}}\frac{\big{\|}\mu^{-1}{\mathcal{D}}\widetilde{f}\big{\|}_{{\cal H}_{0}}\left\|{\mathcal{D}}v_{N}\right\|_{{\cal H}_{0}}}{\left\|v_{N}\right\|_{{\cal H}}}$
$\leq C_{{\rm dis},N,2}\left\|\mu^{-1}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\big{\|}\widetilde{f}\big{\|}_{{\cal H}}.$
By (5.3), (6.9), and (6.7), $\left\|u_{N}\right\|_{{\cal H}}=\left\|\mathbf{U}\right\|_{\mathsf{D}}$ and $\big{\|}\widetilde{f}\big{\|}_{{\cal H}}=\left\|\mathbf{F}\right\|_{\mathsf{D}}$, and the first bound in (6.5) follows. Similarly, by the bound (6.1a) involving $C_{{\rm dis},N,2}$, the definitions of $w_{N}$ (6.8), and the fact that $\left\|\cdot\right\|_{{\cal H}_{0}}\leq\left\|\cdot\right\|_{\cal H}$,
$\left\|w_{N}\right\|_{{\cal H}}\leq C_{{\rm dis},N,2}\sup_{v_{N}\in{\cal H}_{N}}\frac{\big{|}a_{2}(w_{N},v_{N})\big{|}}{\left\|v_{N}\right\|_{{\cal H}}}$ $\leq C_{{\rm dis},N,2}\sup_{v_{N}\in{\cal H}_{N}}\frac{\big{\|}\epsilon\widetilde{f}\big{\|}_{{\cal H}_{0}}\left\|v\right\|_{{\cal H}_{0}}}{\left\|v\right\|_{{\cal H}}}$
$\leq C_{{\rm dis},N,2}\left\|\epsilon\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\big{\|}\widetilde{f}\big{\|}_{{\cal H}_{0}}.$ (6.10)
The second bound in (6.5) then follows since, by (5.3), (6.9), and (6.7), $\left\|w_{N}\right\|_{{\cal H}}=\left\|\mathbf{W}\right\|_{\mathsf{D}}$ and $\|\widetilde{f}\|_{{\cal H}_{0}}\leq\|\widetilde{f}\|_{{\cal H}}=\left\|\mathbf{F}\right\|_{\mathsf{D}}$.
Finally, the bound (6.6) follows from (6.10) by using $m_{-}\|\mathbf{W}\|_{2}\leq\|w_{N}\|_{{\cal H}_{0}}\leq\|w_{N}\|_{{\cal H}}$ and $\|\widetilde{f}\|_{{\cal H}_{0}}\leq m_{+}\|\mathbf{F}\|_{2}$ (with both of these bounds following from (5.4)). ∎
6.2 Norm of inverse solution operator under perturbation
Lemma 6.2 (Norm of inverse solution operator under perturbation).
Suppose that the assumptions of §2 hold. If ${\mathcal{A}}_{1}:{\cal H}\to{\cal H}^{*}$ is invertible and
$\Big{(}\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big{)}\|\mathcal{A}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}\leq 1/2,$ (6.11)
then
$\|\mathcal{A}_{2}^{-1}\|_{{\cal H}^{*}\to{\cal H}}\leq 2\|\mathcal{A}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}.$ (6.12)
Proof.
Given $F\in{\cal H}^{*}$, let $u_{2}\in{\cal H}$ be the solution to $a_{2}(u_{2},v)=F(v)$ for all $v\in{\cal H}$. By the definition of $a_{\ell}(\cdot,\cdot)$ (2.1),
$a_{1}(u_{2},v)=F(v)+\big{(}(\mu^{-1}_{1}-\mu^{-1}_{2}){\mathcal{D}}u_{2},{\mathcal{D}}v\big{)}_{{\cal H}_{0}}-\big{(}(\epsilon_{1}-\epsilon_{2})u_{2},v\big{)}_{{\cal H}_{0}}.$
By the definition of $\|\mathcal{A}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}$, the fact that $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$, and the bound (6.11),
$\left\|u_{2}\right\|_{{\cal H}}$ $\leq\|\mathcal{A}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}\Big{(}\left\|F\right\|_{{\cal H}^{*}}+\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\left\|u_{2}\right\|_{{\cal H}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\left\|u_{2}\right\|_{{\cal H}_{0}}\Big{)}$
$\leq\|\mathcal{A}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}\left\|F\right\|_{{\cal H}^{*}}+\frac{1}{2}\left\|u_{2}\right\|_{{\cal H}},$
and the result (6.12) follows. ∎
Remark 6.3 (Proof of the bound (5.14) under the condition (5.13)).
The condition (5.13) is the same as (6.11), and thus (6.12) holds. Since
$I-{\mathcal{A}}_{2}^{-1}{\mathcal{A}}_{1}={\mathcal{A}}_{2}^{-1}({\mathcal{A}}_{2}-{\mathcal{A}}_{1})\quad\text{ and }\quad I-{\mathcal{A}}_{1}{\mathcal{A}}_{2}^{-1}=({\mathcal{A}}_{2}-{\mathcal{A}}_{1}){\mathcal{A}}_{2}^{-1}.$
the bound (5.14) follows from (6.12) and the bound
$\left\|{\mathcal{A}}_{2}-{\mathcal{A}}_{1}\right\|_{{\cal H}\to{\cal H}^{*}}\leq\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}};$
this last bound holds since, by the definition of $a_{\ell}(\cdot,\cdot)$ (2.1),
$\big{|}a_{1}(u,v)-a_{2}(u,v)\big{|}\leq\Big{(}\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big{)}\left\|u\right\|_{\cal H}\left\|v\right\|_{{\cal H}}$
for all $u,v\in{\cal H}$ (since $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$).
Lemma 6.4 (Norm of discrete solution operator under perturbation).
Suppose that the assumptions of §2 hold. Suppose that, with $0<C_{{\rm dis},N,1}<\infty$,
$\inf_{u_{N}\in{\cal H}_{N}\setminus\{0\}}\sup_{v_{N}\in{\cal H}_{N}\setminus\{0\}}\frac{\big{|}a_{1}(u_{N},v_{N})\big{|}}{\left\|u_{N}\right\|_{{\cal H}}\left\|v_{H}\right\|_{{\cal H}}}\geq\frac{1}{C_{{\rm dis},N,1}},$ (6.13a)
[…]
and
$\Big{(}\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big{)}C_{{\rm dis},N,1}\leq 1/2.$ (6.14)
Then (6.1a) and (6.1b) hold with
$C_{{\rm dis},N,2}=2C_{{\rm dis},N,1}.$
Proof.
Let ${\mathcal{A}}_{N,\ell}:{\cal H}_{N}\to({\cal H}_{N})^{*}$, $\ell=1,2,$ be the operators associated to the sesquilinear forms $a_{\ell}(\cdot,\cdot):{\cal H}_{N}\times{\cal H}_{N}\to\mathbb{C}$, $\ell=1,2$. By Theorem 4.1, the condition (6.13) is equivalent to the statement that
$\|{\mathcal{A}}_{N,1}^{-1}\|_{({\cal H}_{N})^{*}\to{\cal H}_{N}}\leq C_{{\rm dis},N,1}.$ (6.15)
The condition (6.14) then implies that (6.11) holds with $\|\mathcal{A}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}$ replaced by $\|{\mathcal{A}}_{N,1}^{-1}\|_{({\cal H}_{N})^{*}\to{\cal H}_{N}}$. The arguments in the proof of Lemma 6.2 then show that
$\big{\|}{\mathcal{A}}_{N,2}^{-1}\big{\|}_{({\cal H}_{N})^{*}\to{\cal H}_{N}}\leq 2\big{\|}{\mathcal{A}}_{N,1}^{-1}\big{\|}_{({\cal H}_{N})^{*}\to{\cal H}_{N}}.$
Then, by Theorem 4.1 applied to ${\mathcal{A}}_{N,2}$ and (6.15),
$\inf_{u_{N}\in{\cal H}_{N}\setminus\{0\}}\sup_{v_{N}\in{\cal H}_{N}\setminus\{0\}}\frac{\big{|}a_{2}(u_{N},v_{N})\big{|}}{\left\|u_{N}\right\|_{{\cal H}}\left\|v_{H}\right\|_{{\cal H}}}\geq\frac{1}{2\big{\|}{\mathcal{A}}_{N,1}^{-1}\big{\|}_{({\cal H}_{N})^{*}\to{\cal H}_{N}}}\geq\frac{1}{2C_{{\rm dis},N,1}},$
and the result follows. ∎
6.3 Proof of Theorem 5.2
The assumption (5.5) implies that (6.13) holds with $C_{{\rm dis},N,1}=C_{1}\|{\mathcal{A}}^{-1}_{1}\|_{{\cal H}^{*}\to{\cal H}}$. The condition (5.6) is then (6.14) and thus Lemma 6.4 implies that (6.1a) and (6.1b) hold with
$C_{{\rm dis},N,2}=2C_{1}\|{\mathcal{A}}^{-1}_{1}\|_{{\cal H}^{*}\to{\cal H}}.$
With this value of $C_{{\rm dis},N,2}$, the bounds (6.2) and (6.3) from Lemma 6.1 then become the results (5.7) and (5.8).
6.4 Proofs of Lemma 4.2 and 5.9 (about equivalence of the norms of the continuous and discrete inverses)
We first prove Lemma 4.2, which consists of proving the bounds (4.2) and (4.3).
Proof of (4.2).
We highlight that this argument goes back to at least [1, Lemma 3.4], but since it is short, and crucial to the proof of Lemma 5.9, we include it for completeness. The first inequality in (4.2) is an immediate consequence of the fact that $\|\cdot\|_{{\cal H}^{*}}\leq\|\cdot\|_{{\cal H}_{0}}$ (which follows from the definition of $\|\cdot\|_{{\cal H}^{*}}$ and $\|\cdot\|_{{\cal H}_{0}}\leq\|\cdot\|_{{\cal H}}$). Let $a^{+}_{\ell}(u,v):=a_{\ell}(u,v)+C_{\rm G2,\ell}(u,v)_{{\cal H}_{0}}$ and observe that $a^{+}_{\ell}(\cdot,\cdot)$ is coercive by (2.2). To prove the second inequality in (4.2) it is sufficient to prove that, given $F\in{\cal H}^{*}$, the solution of $a_{\ell}(u,v)=F(v)$ for all $v\in{\cal H}$ satisfies
$\left\|u\right\|_{\cal H}\leq(C_{\rm G1})^{-1}\big{(}1+C_{\rm G2,\ell}\big{\|}{\mathcal{A}}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}\big{)}\left\|F\right\|_{{\cal H}^{*}}.$ (6.16)
Let $u^{\pm}$ by the solutions to
$a^{+}_{\ell}(u^{+},v)=F(v)\quad\text{ and }\quad a_{\ell}(u^{-},v)=C_{\rm G2}(u^{+},v)_{{\cal H}_{0}}\quad\text{ for all }v\in{\cal H};$
these definitions imply that $u=u^{+}+u^{-}$. By (2.2) and the Lax–Milgram lemma,
$\big{\|}u^{+}\big{\|}_{{\cal H}}\leq(C_{\rm G1})^{-1}\left\|F\right\|_{{\cal H}^{*}}.$ (6.17)
By the definition of $\|{\mathcal{A}}_{\ell}^{-1}\|_{{\cal H}_{0}\to{\cal H}}$ and the bound (6.17) on $u^{+}$,
$\big{\|}u^{-}\big{\|}_{{\cal H}}\leq\|{\mathcal{A}}_{\ell}^{-1}\|_{{\cal H}_{0}\to{\cal H}}C_{\rm G2,\ell}\big{\|}u^{+}\big{\|}_{{\cal H}_{0}}\leq\|{\mathcal{A}}_{\ell}^{-1}\|_{{\cal H}_{0}\to{\cal H}}C_{\rm G2,\ell}(C_{\rm G1})^{-1}\left\|F\right\|_{{\cal H}^{*}};$ (6.18)
the bound (6.16) – and hence also the second bound in (4.2) – then follows.
Proof of (4.3).
The first inequality in (4.3) is an immediate consequence of the fact that $\|\cdot\|_{{\cal H}_{0}}\leq\|\cdot\|_{{\cal H}}$. To prove the second inequality, observe that the Gårding-type inequality (2.2) implies that, given $f\in{\cal H}_{0}$,
$C_{\rm G1,\ell}\big{\|}\mathcal{A}^{-1}_{\ell}f\big{\|}_{{\cal H}}^{2}$ $\leq C_{\rm G2,\ell}\big{\|}\mathcal{A}^{-1}_{\ell}f\big{\|}^{2}_{{\cal H}_{0}}+\big{|}\langle\mathcal{A}^{-1}_{\ell}f,f\rangle\big{|},$
$\leq C_{\rm G2,\ell}\big{\|}\mathcal{A}^{-1}_{\ell}f\big{\|}^{2}_{{\cal H}_{0}}+\big{\|}\mathcal{A}^{-1}_{\ell}f\big{\|}_{{\cal H}_{0}}\|f\|_{{\cal H}_{0}}.$
Therefore,
$C_{\rm G1,\ell}\big{\|}\mathcal{A}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}}^{2}\leq C_{\rm G2,\ell}\big{\|}\mathcal{A}^{-1}_{\ell}\big{\|}^{2}_{{\cal H}_{0}\to{\cal H}_{0}}+\big{\|}\mathcal{A}^{-1}_{\ell}\big{\|}_{{\cal H}_{0}\to{\cal H}_{0}},$
and the second bound in (4.3) follows.
Finally, to prove Lemma 5.9 we observe the argument above proving the second bound in (4.2) remains unchanged when $F\in{\cal H}^{*}$ is replaced by $F\in({\cal H}_{N})^{*}$; indeed, the Lax–Milgram lemma proves that
$\big{\|}u^{+}_{N}\big{\|}_{{\cal H}}\leq(C_{\rm G1,\ell})^{-1}\left\|F\right\|_{({\cal H}_{N})^{*}}.$
Furthermore, since $u_{N}^{+}\subset{\cal H}_{N}\subset{\cal H}\subset{\cal H}_{0}$, (5.15) implies that
$\big{\|}u^{-}_{N}\big{\|}_{{\cal H}}\leq C_{\rm sol,\ell}C_{\rm G2,\ell}\big{\|}u^{+}\big{\|}_{{\cal H}_{0}}\leq C_{\rm sol,\ell}C_{\rm G2,\ell}(C_{\rm G1,\ell})^{-1}\left\|F\right\|_{({\cal H}_{N})^{*}}$
(compare to (6.18)). The result of Lemma 5.9 therefore follows.
Acknowledgements
The author thanks Qiya Hu (Academy of Mathematics and Systems Science, Chinese Academy of Sciences) for asking him the question of whether the results of [16, 20] extend to the time-harmonic Maxwell equations, and Théophile Chaumont-Frelet (INRIA, Lille) for discussions about the discrete inf-sup constant of the Helmholtz sesquilinear form.