### v2

# Preconditioning FEM discretisations of the high-frequency Helmholtz and Maxwell equations by either perturbing the coefficients or adding absorption

Abstract
This paper investigates the following question: given a Galerkin matrix corresponding to a finite-element discretisation of either the Helmholtz or time-harmonic Maxwell equations with variable coefficients, suppose that the coefficients of the underlying PDE are perturbed; how good an approximate inverse (i.e., preconditioner) is the resulting Galerkin matrix to the original Galerkin matrix? An important special case is when the perturbation consists of adding absorption (in the spirit of “shifted Laplacian preconditioning”). The results of this paper improve the Helmholtz results in [18, 22] and extend these results to the time-harmonic Maxwell equations, confirming a conjecture in the recent preprint [29].
1 Introduction
1.1 Context and motivation
The design of good preconditioners for the linear systems arising from finite-element discretisations of high-frequency time-harmonic wave problems – such as the Helmholtz or time-harmonic Maxwell equations – is a longstanding open problem; see, e.g., the review articles [11, 14, 20, 19] and the references therein. This paper considers the theory of the following two, related, questions involving preconditioning these linear systems.
Preconditioning using absorption.
The idea of preconditioning discretisations of the Helmholtz equation with (approximations of) discretisations of the same Helmholtz problem with added absorption was introduced in [13, 12]. Seeking to rigorously understand this method, the paper [18] proved bounds on ${\mathsf{I}}-\mathsf{A}_{\varepsilon}^{-1}\mathsf{A}$, where $\mathsf{A}$ is the Helmholtz Galerkin matrix, and $\mathsf{A}_{\varepsilon}$ is the Helmholtz Galerkin matrix with $k^{2}\mapsto k^{2}+{\rm i}\varepsilon$, where $k$ is the (real and large) wavenumber. [18] proved that, for a particular nontrapping Helmholtz problem, $\|{\mathsf{I}}-\mathsf{A}_{\varepsilon}^{-1}\mathsf{A}\|_{2}\leq C\varepsilon/k$. An important question is then whether the action of $\mathsf{A}_{\varepsilon}^{-1}$ can be (provably) efficiently approximated when $\varepsilon\sim k$. This question was investigated for one-level domain-decomposition methods in [23], and the recent work [27] proved that certain hybrid two-level Schwarz preconditioners (with problem-adapted basis functions in the coarse space) formed with absorption $\varepsilon\sim k$ give a $k$-independent number of GMRES iterations when applied to the original Helmholtz problem.
“Nearby preconditioning”.
The results of [18] were generalised in [22] to bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$, where $\mathsf{A}_{\ell}$, $\ell=1,2$, are the Galerkin matrices of $k^{-2}\nabla\cdot(A_{\ell}\nabla u_{\ell})+n_{\ell}u_{\ell}=f$. The motivation for studying this situation comes from uncertainty quantification (UQ): calculating quantities of interest for the solution of the Helmholtz equation $k^{-2}\nabla\cdot(A\nabla u)+nu=f$ with random coefficients requires the solution of many deterministic Helmholtz problems, each with different coefficients $A$ and $n$. Bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$ then indicate to what extent a previously-calculated inverse of one of the Galerkin matrices can be used as a preconditioner for other Galerkin matrices; see [36, §4.6] and the recent work [39] for UQ algorithms using this idea.
Content and motivation for the present paper.
The present paper
  • proves the Helmholtz results of [18, 22] under less-restrictive assumptions than in [18, 22], and
  • generalises these results to the time-harmonic Maxwell equations.
Our main motivation is the recent work [29]: this paper gives the first frequency-explicit analysis of a two-level domain-decomposition method for the high-frequency time-harmonic Maxwell equations, under the assumption that the Helmholtz results of [18, 22] hold also for the time-harmonic Maxwell equations; the present paper shows that this assumption indeed holds.
1.2 Informal statement of the main result
Set-up at the continuous level.
[…]
$k^{-2}\nabla\cdot(\mu_{\ell}^{-1}\nabla)-\epsilon_{\ell}$ (1.1)
or the time-harmonic Maxwell operator
$k^{-2}{\rm curl}\big(\mu^{-1}_{\ell}{\rm curl}\big)-\epsilon_{\ell}$ (1.2)
[…]
Specific Helmholtz and Maxwell problems to which the main result applies are described in Lemmas 2.2 and 2.4 below; we highlight that these include Helmholtz and Maxwell problems where the radiation condition is approximated by either a perfectly-matched layer (PML) or an impedance boundary condition.
Set-up at the discrete level.
Given a finite-dimensional subspace ${\cal H}_{N}\subset{\cal H}$ (with $N={\rm dim}({\cal H}_{N})$), let $\mathsf{A}_{\ell},\ell=1,2,$ be the Galerkin matrices associated to $a_{\ell}(\cdot,\cdot)$. Let the matrix $\mathsf{D}$ be such that
$\left\|v_{N}\right\|_{{\cal H}}^{2}=\left\|\mathbf{V}\right\|_{\mathsf{D}}^{2}\quad\text{ for }v_{N}=\sum_{j=1}^{N}V_{j}\phi_{j}\,\text{ with }\,V_{j}\in\mathbb{C}^{N};$ (1.3)
[…]
Informal statement of the main result (Theorem 4.2 below).
The main result gives sufficient conditions under which ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$ (measured in either $\|\cdot\|_{\mathsf{D}}$ or the Euclidean norm $\|\cdot\|_{2}$) is controlled by $\|\mu_{1}-\mu_{2}\|_{L^{\infty}}$ and $\|\epsilon_{1}-\epsilon_{2}\|_{L^{\infty}}$ multiplied by the continuous inf-sup constant of $a_{1}(\cdot,\cdot)$.
In more detail, suppose that ${\cal H}_{N}$ is sufficiently large so that the discrete inf-sup constant of $a_{1}(\cdot,\cdot)$ is bounded by a constant (independent of $N$) multiplied by the continuous inf-sup constant of $a_{1}(\cdot,\cdot)$; denote the latter by $\|{\mathcal{A}}^{-1}_{1}\|_{{\cal H}^{*}\to{\cal H}}$ (see §3 below for an explanation of this notation). If
$\Big(\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{L^{\infty}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{L^{\infty}}\Big)C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}\leq 1/2$
then $\mathsf{A}_{2}^{-1}$ exists and
$\max\Big\{\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{\mathsf{D}},\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{\mathsf{D}^{-1}}\Big\}$
$\hskip 85.35826pt\leq 2\Big(\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{L^{\infty}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{L^{\infty}}\Big)C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}.$ (1.4)
Furthermore, if $\mu_{1}=\mu_{2}$, then there exists $C_{2}$ (independent of $N$ and $k$) such that
$\max\Big\{\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{2},\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{2}\Big\}\leq C_{2}\left\|\epsilon_{1}-\epsilon_{2}\right\|_{L^{\infty}}C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}},$ (1.5)
where $\|\cdot\|_{2}$ denotes the Euclidean norm on matrices (induced by the Euclidean norm on vectors).
1.3 Discussion
1.3.1 The main result specialised to preconditioning with absorption
In the set-up of §1.2, let $\mu_{2}=\mu_{1}$ and $\epsilon_{2}=(1+{\rm i}\alpha)\epsilon_{1}$ with $\alpha>0$ (i.e., $a_{2}(\cdot,\cdot)$ is formed by addition absorption to $a_{1}(\cdot,\cdot)$). If
$\alpha C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}\leq 1/2$
(i.e., the absorption is sufficiently small, depending on the inf-sup constant of $a_{1}(\cdot,\cdot)$), then $\mathsf{A}_{2}^{-1}$ exists,
$\max\Big\{\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{\mathsf{D}},\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{\mathsf{D}^{-1}}\Big\}\leq 2\alpha C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}$ (1.6)
and
$\max\Big\{\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{2},\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{2}\Big\}\leq C_{2}\alpha C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}.$ (1.7)
1.3.2 Recovering the Helmholtz results of [18] about preconditioning with absorption
[18] considers preconditioning the Helmholtz equation with absorption by letting $k^{2}\mapsto k^{2}+{\rm i}\varepsilon$. In the set-up above this corresponds to setting $\alpha:=\varepsilon/k^{2}$. By Remark 3.3 below, $\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}\sim k$ when the Helmholtz problem is nontrapping. The right-hand side of (1.7) is then proportional to $\varepsilon/k$, which is the result proved in [18, Theorem 1.4] for the specific case of the Helmholtz interior impedance problem.
1.3.3 The condition that the discrete inf-sup constant is bounded by the continuous inf-sup constant
As stated in §1.2, the main result holds under the assumption that the dimension of the finite-dimensional approximation space is large enough so that the discrete inf-sup constant is bounded by a constant multiple of the continuous inf-sup constant.
This condition on the discrete inf-sup constant has been proven to hold for Helmholtz problems in the so-called asymptotic regime, i.e., when the finite-dimensional space is sufficiently large so that the Galerkin solution is quasi-optimal; see [32, Theorem 4.2] and the discussion in Remark 4.12 below. Recall that, for the $h$-version FEM applied to Helmholtz non-trapping problems, the asymptotic regime is when $(kh)^{p}k$ is sufficiently small [32, 6].
In §4.3 we prove a new result showing that the discrete inf-sup constant is bounded by a constant multiple of the continuous inf-sup constant in the preasymptotic regime when, for the $h$-version FEM applied to Helmholtz non-trapping problems, $(kh)^{2p}k$ is sufficiently small [17].
Given a Helmholtz or Maxwell problem of interest, the “recipe” to use the main result of this paper is therefore the following:
  1. 1.
    Show that the problem satisfies Assumption 2.1 below. Note that Lemmas 2.2 and 2.4 show that many commonly-studied Helmholtz and Maxwell problems satisfy this assumption.
  2. 2.
    Determine sufficient conditions for the discrete inf-sup constant to be bounded by a constant multiple of the continuous inf-sup constant – these conditions will involve regularity assumptions on the domain and coefficients, as well as conditions on the dimension of the approximation space. Section 4.3 below details these conditions (with references) for the $h$-FEM applied to the Helmholtz and Maxwell problems in Lemmas 2.2 and 2.4.
1.3.4 How the results of this paper improve the Helmholtz results of [18, 21]
Remark 4.7 below discusses in detail how the results of this paper improve the Helmholtz results of [18, 21]. We highlight here that the main way (aside from extending the results to the time-harmonic Maxwell equations) is in proving that bounds (1.4)-(1.7) hold if the discrete inf-sup constant is bounded by a constant multiple of the continuous inf-sup constant – the analyses in [18, 21] require more restrictive conditions on the finite-dimensional space.
1.3.5 Numerical experiments investigating the sharpness of the bounds (1.4), (1.5), (1.6), and (1.7)
[18, §5] and [22, §2.2.2] contain numerical experiments investigating the sharpness of the bounds (1.7) and (1.4)/(1.5) respectively. We highlight that these experiments consider piecewise-linear FEM discretisations in the preasymptotic regime, i.e., $(kh)^{2p}k$ with $p=1$; these FEM discretisations were not covered by the theory in [18, 22] (see the discussion in §1.3.3 and in Remark 4.7 below) but are covered by the results of the present paper.
1.3.6 The use of the bound (1.6) in the DD analyses of [27] and [29]
The Helmholtz two-level DD analysis of [27] assumes that the bound (1.6) holds for the Helmholtz interior impedance problem (see the assumptions of [27, Corollary 6.1]). Similarly, the Maxwell two-level DD analysis of [29] assumes that the bound (1.6) holds for the Maxwell interior impedance problem (see the discussion on [29, Page 2]). The main result of this paper, Theorem 4.2, coupled with results about the discrete inf-sup constant in §4.3 below then give sufficient conditions on the finite-element spaces for these assumptions to be satisfied.
Outline of the rest of the paper.
§2 states the abstract assumptions under which the main results are proved and describes Helmholtz and Maxwell problems that are covered by these abstract assumptions. §3 recaps properties of the inf-sup constant of a sesquilinear form and its $k$-dependence for Helmholtz and Maxwell. §4 states the main results, and §5 proves them.
2 The class of Helmholtz and Maxwell problems considered
The main result is proved under the following abstract assumptions.
Assumption 2.1 (Abstract assumptions).
[…]
$a_{\ell}(u,v):=\big(\mu_{\ell}^{-1}{\mathcal{D}}u,{\mathcal{D}}v\big)_{{\cal H}_{0}}+b(u,v)-\big(\epsilon_{\ell}u,v\big)_{{\cal H}_{0}},\quad\ell=1,2,$ (2.1)
where $\mu_{\ell}^{-1}:{\cal H}_{0}\to{\cal H}_{0}$ and $\epsilon_{\ell}:{\cal H}_{0}\to{\cal H}_{0}$ are bounded linear operators for $\ell=1,2$. Finally for $\ell=1,2,$ there exists $C_{\rm G1,\ell},C_{\rm G2,\ell}>0$ such that
$\big|a_{\ell}(v,v)+C_{\rm G2,\ell}\left\|v\right\|^{2}_{{\cal H}_{0}}\big|\geq C_{\rm G1,\ell}\left\|v\right\|^{2}_{{\cal H}}\quad\text{ for all }v\in{\cal H}.$ (2.2)
We make three immediate remarks.
  • These assumptions imply that $a_{\ell}(\cdot,\cdot)$ is continuous; i.e., there exist $C_{\rm cont,\ell}>0$ such that
    $\big|a_{\ell}(u,v)\big|\leq C_{\rm cont,\ell}\left\|u\right\|_{{\cal H}}\left\|v\right\|_{{\cal H}}\quad\text{ for all }u,v\in{\cal H}.$
[…]
    $\Re a_{\ell}(v,v)\geq C_{\rm G1,\ell}\left\|v\right\|^{2}_{{\cal H}}-C_{\rm G2,\ell}\left\|v\right\|^{2}_{{\cal H}_{0}}\quad\text{ for all }v\in{\cal H},$ (2.3)
    then (2.2) holds (since $|z|\geq\Re z$ for all $z\in\mathbb{C}$).
  • If $a_{\ell}(\cdot,\cdot)$ satisfies these assumptions, then so does
    $a_{\ell}^{*}(u,v):=\overline{a_{\ell}(v,u)}=\big(\mu_{\ell}^{*}{\mathcal{D}}u,{\mathcal{D}}v\big)_{{\cal H}_{0}}+\overline{b(v,u)}-\big(\epsilon_{\ell}^{*}u,v\big)_{{\cal H}_{0}},$
    with the same $C_{\rm G1,\ell},C_{\rm G2,\ell}$ (this is because $|a_{\ell}(v,v)+C_{\rm G2,\ell}\left\|v\right\|^{2}_{{\cal H}_{0}}|=\big|\overline{a_{\ell}(v,v)}+C_{\rm G2,\ell}\left\|v\right\|^{2}_{{\cal H}_{0}}\big|$).
We now give examples of Helmholtz and Maxwell problems satisfying Assumption 2.1.
Lemma 2.2 (Helmholtz problems satisfying Assumption 2.1).
Let $\Omega$ be a bounded Lipschitz open set, ${\cal H}_{0}=L^{2}(\Omega)$ (scalar valued), ${\mathcal{D}}:=k^{-1}\nabla$, and
$\left\|v\right\|^{2}_{{\cal H}}:=\left\|k^{-1}\nabla v\right\|^{2}_{L^{2}(\Omega)}+\left\|v\right\|^{2}_{L^{2}(\Omega)}.$ (2.4)
For $\ell=1,2,$ let $\mu_{\ell}^{-1}$ be bounded, symmetric, matrix functions on $\Omega$ with ${\rm essinf}_{\Omega}\Re\mu_{\ell}^{-1}>0$ (in the sense of quadratic forms) and $\epsilon_{\ell}$ be bounded, scalar-valued functions on $\Omega$.
If one of the following three points holds, then Assumption 2.1 holds.
[…]
and $b(\cdot,\cdot)=0$.
(ii) With $\Gamma_{\rm imp}$ a non-empty connected component of $\partial\Omega$, and $\Gamma_{\rm D}$ a (possibly empty) connected component of $\partial\Omega$ that is not equal to $\Gamma_{\rm imp}$,
${\cal H}:=\big\{v\in H^{1}(\Omega)\,:\,v=0\text{ on }\Gamma_{\rm D}\big\},$
and
$b(u,v):=-{\rm i}k^{-1}\big(\theta u,v\big)_{L^{2}(\Gamma_{\rm imp})}$
for some real-valued $\theta\in L^{\infty}(\Gamma_{\rm imp})$ such that ${\rm essinf}_{\Gamma_{\rm imp}}\theta>0$.
[…]
$b(u,v):=-{\rm i}k^{-1}\big({\rm DtN}_{k}u,v\big)_{L^{2}(\Gamma_{\rm imp})},$
where ${\rm DtN}_{k}$ is the exact Dirichlet-to-Neumann map for the Helmholtz equation in the exterior of $\{x:|x|\leq R\}$; see, e.g., [2, Equations 3.5 and 3.6] for explicit expressions in terms of Fourier series (in 2-d) or spherical harmonics (in 3-d) and Bessel and Hankel functions.
Remark 2.3 (The Helmholtz problems covered in Lemma 2.2).
The fact that ${\mathcal{D}}:=k^{-1}\nabla$ implies that the sesquilinear form (2.1) corresponds to the Helmholtz operator (1.1) Parts (i), (ii), and (iii) of Lemma 2.2 then cover three common ways of approximating the Sommerfeld radiation condition (i.e., the fact the scattered wave travels “outward” from the scatterer; see, e.g., [7, §3.2]).
Indeed, in Part (i), zero Dirichlet boundary conditions are imposed on at least one connected component of $\partial\Omega$, with then zero Neumann boundary conditions imposed on the rest. The requirements on $\mu_{\ell}$ and $\epsilon_{\ell}$ in Lemma 2.2 are then satisfied by the coefficients coming from a radial perfectly-matched layer (PML) by, e.g., [16, Lemma 2.3].
In Part (ii), the sesquilinear form corresponds to the Helmholtz operator (1.1) with the impedance boundary condition
$k^{-1}\partial_{n,\mu^{-1}}u-{\rm i}\theta u=g_{\rm imp}\quad\text{ on }\Gamma_{\rm imp}$
$u=0$ on $\Gamma_{\rm D}$, and $k^{-1}\partial_{n,\mu^{-1}}u=g_{\rm N}$ on $\partial\Omega\setminus(\Gamma_{\rm D}\cup\Gamma_{\rm imp})$ (where the data $g_{\rm imp}$ and $g_{\rm N}$ depend on the given right-hand side in the variational problem). Here $\partial_{n,\mu^{-1}}$ is the conormal derivative; recall that this is such that $\partial_{n,\mu^{-1}}u:=n\cdot\mu^{-1}\nabla u$ when $u\in H^{2}(\Omega)$, where $n$ is the outward-pointing unit normal vector to $\partial\Omega$ (see, e.g., [31, Lemmas 4.2 and 4.3]).
In Part (iii), the sesquilinear form corresponds to the Helmholtz operator (1.1) with the exact Dirichlet-to-Neumann map imposed on the boundary of a ball.
Proof of Lemma 2.2.
We need to check that (a) $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$, (b) $b(\cdot,\cdot)$ is continuous, and (c) the inequality (2.2) holds.
Regarding (a): this follows immediately from the definition of the norm (2.4) and the fact that ${\mathcal{D}}:=k^{-1}\nabla$.
Regarding (b): for Part (i), $b(\cdot,\cdot)=0$ and so is automatically continuous. For Part (ii), continuity of $b(\cdot,\cdot)$ follows from the standard multiplicative trace inequality (see, e.g., [24, Theorem 1.5.1.10, last formula on p. 41], [1, Theorem 1.6.6]). For Part (iii), continuity of $b(\cdot,\cdot)$ follows from, e.g., [32, Lemma 3.3, Part 1].
Regarding (c): for Part (i), the inequality (2.2) follows immediately since $b(\cdot,\cdot)=0$. For Part (ii), $a(\cdot,\cdot)$ satisfies the Gårding inequality (2.3) (with the term on $\Gamma_{\rm imp}$ playing no role because of the real part on the left-hand side of (2.3)), and thus (2.2) holds. For Part (iii), recall that $\Re(-{\rm i}{\rm DtN}_{k})\geq 0$ by, e.g., [2, Corollary 3.1] or [32, Lemma 3.3, Part 2], so that $a(\cdot,\cdot)$ satisfies the Gårding inequality (2.3). ∎
Lemma 2.4 (Maxwell problems satisfying Assumption 2.1).
Let $\Omega$ be a bounded Lipschitz open set, ${\cal H}_{0}=L^{2}(\Omega)$ (vector valued), and ${\mathcal{D}}:=k^{-1}{\rm curl}$.
For $\ell=1,2,$ let $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ be bounded, symmetric, matrix functions on $\Omega$ with ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{\rm essinf}_{\Omega}}}\Re\mu_{\ell}^{-1}>0$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{\rm essinf}_{\Omega}}}\Re\epsilon_{\ell}>0$ (in the sense of quadratic forms).
If one of the following two points holds, then Assumption 2.1 holds.
${\cal H}:=H_{0}({\rm curl};\Omega)=\big\{v\in H({\rm curl};\Omega):v_{T}=0\big\}$
where $v_{T}$ is the tangential trace such that $v_{T}:=(n\times v)\times n$ for smooth $v$, where $n$ is the outward-pointing unit normal vector (see, e.g., [34, Equation 3.46 and Theorem 3.31]).
$\left\|v\right\|^{2}_{{\cal H}}:=\left\|k^{-1}{\rm curl}\,v\right\|^{2}_{L^{2}(\Omega)}+\left\|v\right\|^{2}_{L^{2}(\Omega)}$ (2.5)
and $b(\cdot,\cdot)=0$.
(ii) With $\Gamma_{\rm imp}$ a non-empty connected component of $\partial\Omega$,
${\cal H}:=\big\{u\in H^{1}(\Omega)\,:\,\text{ $u_{T}=0$ on $\partial\Omega\setminus\Gamma_{\rm imp}$ and $u_{T}\in L^{2}(\Gamma_{\rm imp})$}\big\},$
$\left\|v\right\|^{2}_{{\cal H}}:=k^{-2}\left\|{\rm curl}\,v\right\|^{2}_{L^{2}(\Omega)}+\left\|v\right\|^{2}_{L^{2}(\Omega)}+k^{-1}\left\|v_{T}\right\|^{2}_{L^{2}(\Gamma_{\rm imp})},$ (2.6)
and
$b(u,v):=-{\rm i}k^{-1}\big(\theta u_{T},v_{T}\big)_{L^{2}(\Gamma_{\rm imp})},$
for some real-valued $\theta\in L^{\infty}(\Gamma_{\rm imp})$ such that ${\rm essinf}_{\Gamma_{\rm imp}}\theta>0$. In addition, $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ are real-valued.
Remark 2.5 (The Maxwell problems covered in Lemma 2.2).
The fact that ${\mathcal{D}}:=k^{-1}{\rm curl}$ implies that the sesquilinear form (2.1) corresponds to the Maxwell operator and the sesquilinear form corresponds to the Maxwell operator (1.2) Parts (i) and (ii) of Lemma 2.2 then cover two common ways of approximating the radiation condition.
Indeed, the requirements on $\mu_{\ell}$ and $\epsilon_{\ell}$ in Part (i) are satisfied by the coefficients coming from a radial PML by, e.g., [4, Appendix A]. In Part (ii), the sesquilinear form corresponds to the Maxwell operator (1.2) with the impedance boundary condition
$k^{-1}\big(\mu_{\ell}^{-1}{\rm curl}\,u\big)\times n-{\rm i}\theta u_{T}=g\quad\text{ on }\Gamma_{\rm imp}$
(where $g$ depends on the given right-hand side) and the PEC boundary condition $u_{T}=0$ on $\Gamma\setminus\Gamma_{\rm imp}$.
Proof of Lemma 2.4.
As in the Helmholtz case, we need to check that (a) $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$, (b) $b(\cdot,\cdot)$ is continuous, and (c) the inequality (2.2) holds.
Regarding (a): this follows immediately from the definitions of the norm (2.5)/(2.6) and the fact that ${\mathcal{D}}:=k^{-1}{\rm curl}$.
Regarding (b): for Part (i), continuity of $b(\cdot,\cdot)$ is immediate (since it is zero). For Part (ii), continuity of $b(\cdot,\cdot)$ follows from the Cauchy–Schwarz inequality and the definition of the norm (2.6).
Regarding (c): for Part (i), the inequality (2.2) follows immediately since $b(\cdot,\cdot)=0$. For Part (ii), since $\mu^{-1}_{\ell}$ and $\epsilon_{\ell}$ are both real (by assumption), $\Im(a(v,v))$ consists only of $\Im(b(v,v))$, which is bounded below by $k^{-1}({\rm essinf}_{\Gamma_{\rm imp}}\theta)\|v\|^{2}_{L^{2}(\Gamma_{\rm imp})}$; the bound (2.2) then follows. ∎
Remark 2.6 (The reason for the particular weighting with $k$).
[…]
3 Recap of properties of the inf-sup constant of $a_{\ell}(\cdot,\cdot)$
In this section, we recap properties of the inf-sup constant of $a_{\ell}(\cdot,\cdot)$, since it appears in the statement of the main result (Theorem 4.2 below).
[…]
Theorem 3.1 (Inf-sup condition equivalent to bound on inverse operator).
The conditions
$\inf_{u\in{\cal H}\setminus\{0\}}\sup_{v\in{\cal H}\setminus\{0\}}\frac{\big|a_{\ell}(u,v)\big|}{\left\|u\right\|_{{\cal H}}\left\|v\right\|_{{\cal H}}}\geq\gamma^{-1}\text{ and for all }v\in{\cal H}\setminus\{0\},\sup_{u\in{\cal H}\setminus\{0\}}|a_{\ell}(u,v)|>0$ (3.1)
and
$\big\|{\mathcal{A}}_{\ell}^{-1}\big\|_{{\cal H}^{*}\to{\cal H}}\leq\gamma$
are equivalent.
References for the proof.
[…]
The following lemma, proved in §5.4 below using the Gårding-type inequality (2.2), shows that all norms of ${\mathcal{A}}_{\ell}^{-1}$ are equivalent, with constants depending only on $C_{\rm G1,\ell}$ and $C_{\rm G2,\ell}$ (and hence independent of $k$).
Lemma 3.2 ($k$-independent equivalence of norms of ${\mathcal{A}}_{\ell}^{-1}$).
If ${\mathcal{A}}_{\ell}^{-1}:{\cal H}^{*}\to{\cal H}$ exists, then
$\big\|\mathcal{A}_{\ell}^{-1}\big\|_{{\cal H}_{0}\to{\cal H}}\leq\big\|\mathcal{A}_{\ell}^{-1}\big\|_{{\cal H}^{*}\to{\cal H}}\leq(C_{\rm G1,\ell})^{-1}\Big(1+C_{\rm G2,\ell}\big\|\mathcal{A}_{\ell}^{-1}\big\|_{{\cal H}_{0}\to{\cal H}}\Big)$ (3.2)
and
$\big\|\mathcal{A}_{\ell}^{-1}\big\|_{{\cal H}_{0}\to{\cal H}_{0}}\leq\big\|\mathcal{A}_{\ell}^{-1}\big\|_{{\cal H}_{0}\to{\cal H}}\leq(C_{\rm G1,\ell})^{-1/2}\big\|\mathcal{A}_{\ell}^{-1}\big\|_{{\cal H}_{0}\to{\cal H}_{0}}\sqrt{C_{\rm G2,\ell}+\big\|\mathcal{A}_{\ell}^{-1}\big\|_{{\cal H}_{0}\to{\cal H}_{0}}^{-1}}.$ (3.3)
Remark 3.3 (The $k$-dependence of $\|{\mathcal{A}}^{-1}_{\ell}\|$ for Helmholtz and Maxwell).
When $\mu_{\ell}$ and $\epsilon_{\ell}$ are both constant multiples of the identity in part of the domain, $\|{\mathcal{A}}^{-1}_{\ell}\|\geq Ck$ – this can be proved by considering data that is a cut-off function multiplied by a plane wave; see, e.g., [2, Lemma 3.10] (for Helmholtz) and [5, §1.4.1], [33, Example 3.4] (for Maxwell).
For the scattering problem, if the scatterer is nontrapping then $\|{\mathcal{A}}^{-1}_{\ell}\|\leq Ck$; this has been proved in a wide variety of Helmholtz cases (see, e.g.,, the literature review in [21]) and for the Maxwell problem with certain nontrapping coefficients in [5] or a nontrapping PEC obstacle in [40, §2]. Under trapping, $\|{\mathcal{A}}^{-1}_{\ell}\|\gg k$; see the results and literature reviews in [28], [3].
For the PML problem, [15, Theorem 1.6] proved that the norm of the solution operator of the Helmholtz radial PML problem is bounded by the norm of the solution operator of the corresponding Helmholtz scattering problem; we expect that the same result holds for the Maxwell PML problem. Indeed, this result was recently proved – up to a factor of $k$ loss – for the case of no scatterer and Cartesian PML in [8, Lemma 10].
4 Statement of the main results
This section states precisely the main results of the paper; the proofs of these results are then given in §5.
4.1 Notation for the Galerkin matrices
Let ${\cal H}_{N}\subset{\cal H}$ be a finite-dimensional space with basis $\{\phi_{j}\}_{j=1}^{N}$. Let
$\mathsf{A}_{\ell}=\mathsf{S}_{\mu^{-1}_{\ell}}+\mathsf{B}-\mathsf{M}_{\epsilon_{\ell}},\quad\ell=1,2,$ (4.1)
where
$(\mathsf{S}_{\mu^{-1}_{\ell}})_{ij}=\big(\mu^{-1}_{\ell}{\mathcal{D}}\phi_{j},{\mathcal{D}}\phi_{i}\big)_{{\cal H}_{0}},\quad(\mathsf{M}_{\epsilon_{\ell}})_{ij}=\big(\epsilon_{\ell}\phi_{j},\phi_{i}\big)_{{\cal H}_{0}},\quad\text{ and }\quad\mathsf{B}_{ij}=b(\phi_{j},\phi_{i}).$ (4.2)
Let the matrix $\mathsf{D}$ be such that the norm relation (1.3) holds. We also use the notation $\|\cdot\|_{\mathsf{D}}$ to denote the induced norm on matrices. Finally, let $\|\cdot\|_{2}$ denote the Euclidean norm on $\mathbb{C}^{N}$, and let $m_{\pm}$ be such that
$m_{-}\|\mathbf{V}\|_{2}\leq\|v_{N}\|_{{\cal H}_{0}}\leq m_{+}\|\mathbf{V}\|_{2}\quad\text{ for all }v_{N}\in{\cal H}_{N}.$ (4.3)
Remark 4.1.
With $\mathsf{M}$ the mass matrix, i.e., $(\mathsf{M})_{ij}=(\phi_{j},\phi_{i})_{{\cal H}_{0}}$, (4.3) is equivalent to the bounds
$m_{-}^{2}\|\mathbf{V}\|_{2}^{2}\leq\big(\mathsf{M}\mathbf{V},\mathbf{V}\big)_{2}\leq m_{+}^{2}\|\mathbf{V}\|_{2}^{2}\quad\text{ for all }\mathbf{V}\in\mathbb{C}^{N};$
i.e., the quantity $(m_{+}/m_{-})^{2}$ (whose square root appears in the bound (4.7) below) is the ratio of the largest to the smallest eigenvalue – i.e., the condition number – of the mass matrix $\mathsf{M}$. For the $h$-FEM with quasi-uniform meshes, the ratio $m_{+}/m_{-}$ is independent of $h$; see, e.g., [18, Equation 4.2], [22, Lemma 5.1], [35, Lemma 4.6].
4.2 The main result: bounds on $\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\|$ and $\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\|$
Theorem 4.2 (Bounds on $\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\|$ and $\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\|$).
Suppose that the assumptions of §2 hold. Suppose that $({\cal H}_{N})_{N=1}^{\infty}$ are such that there exists $C_{1}>0$ such that, for all $N\in\mathbb{Z}^{+}$,
$\inf_{u_{N}\in{\cal H}_{N}\setminus\{0\}}\sup_{v_{N}\in{\cal H}_{N}\setminus\{0\}}\frac{\big|a_{1}(u_{N},v_{N})\big|}{\left\|u_{N}\right\|_{{\cal H}}\left\|v_{H}\right\|_{{\cal H}}}\geq\frac{1}{C_{1}\big\|{\mathcal{A}}_{1}^{-1}\big\|_{{\cal H}^{*}\to{\cal H}}}$ (4.4a)
and
 $\text{ for all }v_{N}\in{\cal H}_{N}\setminus\{0\},\,\sup_{u_{N}\in{\cal H}_{N}\setminus\{0\}}|a_{1}(u_{N},v_{N})|>0$ (4.4b)
(i.e., the discrete inf-sup constant of $a_{1}(\cdot,\cdot)$ is bounded below by a constant times the continuous inf-sup constant (3.1)). If
$\Big(\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big)C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}\leq 1/2$ (4.5)
then $\mathsf{A}_{2}^{-1}$ exists and
$\max\Big\{\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{\mathsf{D}},\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{\mathsf{D}^{-1}}\Big\}$
$\hskip 85.35826pt\leq 2\Big(\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big)C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}.$ (4.6)
Furthermore, if $\mu_{1}=\mu_{2}$, then
$\max\Big\{\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{2},\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{2}\Big\}\leq 2\frac{m_{+}}{m_{-}}\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}.$ (4.7)
We make the following four immediate remarks.
  • (i)
    The assumptions of Theorem 4.2 – by design – involve $a_{1}(\cdot,\cdot)$ and not $a_{2}(\cdot,\cdot)$; i.e., the assumptions are on the problem that we want to solve (involving $\mathsf{A}_{1}$), rather than the problem used for preconditioning (involving $\mathsf{A}_{2}$).
  • (ii)
    Since ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}=\mathsf{A}_{2}^{-1}(\mathsf{A}_{2}-\mathsf{A}_{1})$, the right-hand sides of both (4.6) and (4.7) should be understood as $\|\mathsf{A}_{2}-\mathsf{A}_{1}\|\|\mathsf{A}_{2}^{-1}\|$; i.e., the norm of the perturbation times the norm of the solution operator. The key point is that the conditions (4.4a) and (4.5) mean that $\|\mathsf{A}_{2}^{-1}\|$ is bounded by $\|{\mathcal{A}}_{2}^{-1}\|_{{\cal H}^{*}\to{\cal H}}$, which in turn is bounded by $\|{\mathcal{A}}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}$.
  • (iii)
    §4.3 gives sufficient conditions for (4.4a) to hold (including a new result on this); i.e., conditions under which the discrete inf-sup constant is bounded by a constant times the continuous inf-sup constant.
  • (iv)
    In all our Helmholtz and Maxwell examples, the ${\cal H}_{0}\to{\cal H}_{0}$ norms of the coefficient differences appearing on the right-hand sides of both (4.6) and (4.7) are bounded by their $L^{\infty}$ norms (since $\|\mu v\|_{L^{2}(\Omega)}\leq\|\mu\|_{L^{\infty}(\Omega)}\left\|v\right\|_{L^{2}(\Omega)}$ for all $v\in L^{2}(\Omega)$).
Corollary 4.3 (Theorem 4.2 specialised to preconditioning with absorption).
Suppose that the assumptions of §2 hold with $\mu_{2}=\mu_{1}$ and $\epsilon_{2}=(1+{\rm i}\alpha)\epsilon_{1}$ with $\alpha>0$. Suppose that $({\cal H}_{N})_{N=1}^{\infty}$ are such that there exists $C_{1}>0$ such that (4.4) hold for all $N\in\mathbb{Z}^{+}$. If
$\alpha C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}\leq 1/2$
then $\mathsf{A}_{2}^{-1}$ exists,
$\max\Big\{\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{\mathsf{D}},\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{\mathsf{D}^{-1}}\Big\}\leq 2\alpha C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}$ (4.8)
and
$\max\Big\{\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{2},\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{2}\Big\}\leq 2\frac{m_{+}}{m_{-}}\alpha C_{1}\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}.$ (4.9)
Remark 4.4 (Implications of Theorem 4.2 for preconditioning).
If the right-hand side of (4.6) is $\leq c<1$, then
  • the preconditioned fixed point iteration $\mathbf{x}^{n+1}=\mathbf{x}^{n}+\mathsf{A}_{2}^{-1}({\bf b}-\mathsf{A}_{1}\mathbf{x}^{n})$ for solving $\mathsf{A}_{1}\mathbf{x}={\bf b}$ satisfies
    $\left\|\mathbf{x}-\mathbf{x}^{n}\right\|_{\mathsf{D}}\leq c^{n}\left\|\mathbf{x}-\mathbf{x}^{0}\right\|_{\mathsf{D}},$ (4.10)
  • when GMRES is applied in the $\mathsf{D}$ inner product, the residual $\mathbf{r}^{n}:=\mathsf{A}_{2}^{-1}{\bf b}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\mathbf{x}^{n}$ satisfies
    $\frac{\left\|\mathbf{r}^{n}\right\|_{\mathsf{D}}}{\left\|\mathbf{r}^{0}\right\|_{\mathsf{D}}}\leq\left(\frac{2\sqrt{c}}{(1+c)^{2}}\right)^{n}$ (4.11)
    (as a consequence of the Elman estimate [10, 9]; see [18, Corollary 1.9]/[22, Corollary 5.5]).
Similar statements hold for right-preconditioning and/or the fixed-point iteration/GMRES in the Euclidean inner product (using (4.7) instead of (4.6)).
We highlight that, at least in the case of preconditioning by absorption, the action of $\mathsf{A}_{2}^{-1}$ must be further approximated to provide a practical preconditioner (as discussed in §1.1).
Remark 4.5 (The analogue of Theorem 4.2 at the level of sesquilinear forms).
A byproduct of the proof of Theorem 4.2 below is that if
$\Big(\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big)\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}\leq 1/2$ (4.12)
then ${\mathcal{A}}_{2}^{-1}$ exists and
$\max\Big\{\big\|I-{\mathcal{A}}_{2}^{-1}{\mathcal{A}}_{1}\big\|_{{\cal H}},\big\|I-{\mathcal{A}}_{1}{\mathcal{A}}_{2}^{-1}\big\|_{{\cal H}}\Big\}$
$\hskip 85.35826pt\leq 2\Big(\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big)\big\|{\mathcal{A}}^{-1}_{1}\big\|_{{\cal H}^{*}\to{\cal H}}$ (4.13)
(see Remark 5.3). For the Helmholtz case with $\mu_{1}=\mu_{2}$, [22, Lemma 2.7] exhibits $\epsilon_{1},\epsilon_{2}$ such that the bound (4.13) is sharp in its $k$-dependence (in this example $\|{\mathcal{A}}^{-1}_{1}\|_{{\cal H}^{*}\to{\cal H}}\sim k$).
Remark 4.6 (Measuring the difference in the coefficients in weaker norms).
By using standard finite-element inverse estimates, [22, Lemma 5.3] proved (for the Helmholtz case) bounds on $\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{\mathsf{D}}$, $\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{\mathsf{D}}$, $\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{2}$, and $\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{2}$ with $L^{q}$ norms of both $\mu^{-1}_{1}-\mu^{-1}_{2}$ and $\epsilon_{1}-\epsilon_{2}$ appearing on the right-hand sides. These bounds also hold for the Helmholtz and Maxwell problems in §2; however, we do not include them here since (i) these bounds involves powers of $h^{-1}$ (from the inverse estimates) and are thus worse than (4.6) and (4.7), and (ii) the bounds (4.6) and (4.7) are sufficient for studying preconditioning with absorption (becoming (4.8) and (4.9) in Corollary 4.3) and, in particular, verifying the assumption about this in [29].
Remark 4.7 (Theorem 4.2 improves the Helmholtz results of [18] and [22] in three ways).
The bounds (4.6)/(4.7) and (4.9) are essentially the same as those proved in [22] and [18], respectively. Nevertheless, Theorem 4.2 improves the results of [18, 22] by enlarging the class of Helmholtz problems to which the bounds apply, and weakening the assumptions on the finite-dimensional space. In more detail:
(i) Theorem 4.2 covers all the Helmholtz problems in Lemma 2.2, whereas [18] and [22] only consider the exterior Dirichlet problem with a nontrapping obstacle (with the obstacle allowed to be empty). For this problem, [18] considers truncation of the exterior domain by an impedance boundary condition, and [22] considers truncation by either an impedance boundary condition or the exact Dirichlet-to-Neumann map.
(ii) Theorem 4.2 makes assumptions only about $a_{1}(\cdot,\cdot)$ (i.e., the problem we want to precondition), whereas [18] and [22] make assumptions about both $a_{1}(\cdot,\cdot)$ and $a_{2}(\cdot,\cdot)$; e.g., the mesh threshold in [18, Equation 1.12] is for the problem with absorption, i.e., $a_{2}(\cdot,\cdot)$, and [22, Theorems 2.2 and 2.3] make assumptions on both $a_{1}(\cdot,\cdot)$ and $a_{2}(\cdot,\cdot)$.
(iii) Theorem 4.2 is proved under the discrete inf-sup condition (4.4) (which, by Theorem 3.1, is equivalent to the Galerkin matrix being invertible), whereas the bounds of [18] and [22] are proved under stronger assumptions than a discrete inf-sup condition. Indeed, [18, Equations 1.13 and 1.14] prove bounds analogous to (4.7) under the assumption that, given any data and a sequence of finite-dimensional subspaces, the sequence of Galerkin solutions is quasi-optimal, with constant independent of $k$ – we see in Corollary 4.9 below that this implies that the Galerkin matrix is invertible, with the norm of its inverse bounded by the norm of the PDE solution operator. Furthermore, [22, Lemma 5.3] proves bounds analogous to (4.6) and (4.7) under assumptions weaker than quasi-optimality, but stronger than invertibility of the Galerkin matrix; see [22, Condition 4.2].
4.3 Auxiliary result: conditions under which the discrete inf-sup constant is proportional to the (continuous) inf-sup constant
The main observation in this subsection (which we couldn’t find elsewhere in the literature) is that the argument bounding $\|\mathcal{A}_{\ell}^{-1}\|_{{\cal H}^{*}\to{\cal H}}$ in terms of $\|\mathcal{A}_{\ell}^{-1}\|_{{\cal H}_{0}\to{\cal H}}$ (in Lemma 3.2) also holds at the discrete level, giving the following result.
Lemma 4.8 (Bound on discrete inf-sup constant via bound on Galerkin solution).
Suppose that the assumptions in §2 hold and there exists $C_{\rm sol,\ell}>0$ such that, for all $f\in{\cal H}_{0}$, the solution $u_{N}\in{\cal H}_{N}$ to
$a_{\ell}(u_{N},v_{N})=(f,v_{N})_{{\cal H}_{0}}\quad\text{ for all }v_{N}\in{\cal H}_{N}$
exists and satisfies
$\left\|u_{N}\right\|_{{\cal H}}\leq C_{\rm sol,\ell}\left\|f\right\|_{{\cal H}_{0}}.$ (4.14)
Then
[…]
In other words, the bound (4.14) on the discrete solution for the restricted class of data $f\in{\cal H}_{0}$ (instead of $f\in{\cal H}^{*}$) is sufficient to obtain a bound on the discrete inf-sup constant.
Several analyses of the $h$-FEM for Helmholtz and Maxwell prove the bound (4.14) directly; e.g., this is done in [30, Theorem 4.1] for the Maxwell impedance problem in Part (ii) of Lemma 2.4 when $\Omega$ a $C^{2}$ domain, $\mu_{\ell}=\epsilon_{\ell}=I$, $({\cal H}_{N})_{N=0}^{\infty}$ consists of Nédélec edge elements of the second type, and $(kh)^{2}k$ sufficiently small.
We now give two other ways of proving the bound (4.14), and hence bounding the discrete inf-sup constant. The first of these is in the asymptotic regime; i.e., when the sequence of Galerkin solutions is quasi-optimal – for the $h$-FEM applied to Helmholtz/Maxwell, this is when $(kh)^{p}\|{\mathcal{A}}_{\ell}^{-1}\|$ is sufficiently small; the second of these is in the preasymptotic regime.
Corollary 4.9 (Bound on discrete inf-sup constant via quasi-optimality).
Suppose that the assumptions in §2 hold and there exists $0<C_{\rm qo,\ell}<\infty$ such that the following holds. Suppose that $({\cal H}_{N})_{N=0}^{\infty}$ is a sequence of finite-dimensional subspaces such that, for all $N$, given $f\in{\cal H}_{0}$ the Galerkin solution $u_{N}$ to
$a_{\ell}(u_{N},v_{N})=(f,v_{N})_{{\cal H}_{0}}\quad\text{ for all }v_{N}\in{\cal H}_{N}$
exists, is unique, and satisfies
$\left\|u-u_{N}\right\|_{{\cal H}}\leq C_{\rm qo,\ell}\left\|(I-\Pi_{N})u\right\|_{{\cal H}},$
[…]
Then
$\inf_{u_{h}\in{\cal H}_{h}\setminus\{0\}}\sup_{v_{h}\in{\cal H}_{h}\setminus\{0\}}\frac{\big|a_{\ell}(u_{h},v_{h})\big|}{\left\|u_{h}\right\|_{{\cal H}}\left\|v_{h}\right\|_{{\cal H}}}\geq\frac{1}{(C_{\rm G1,\ell})^{-1}\Big(1+C_{\rm G2,\ell}\big(1+C_{\rm qo,\ell}\big)\big\|\mathcal{A}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}\Big)}.$
Corollary 4.10 (Bound on discrete inf-sup constant in preasymptotic regime).
[…]
Suppose there exist $C_{j},j=1,2,3,$ such that if
$(kh)^{2p}\big\|{\mathcal{A}}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}\leq C_{1}$ (4.15)
then, for all $u\in{\cal H},u_{h}\in{\cal H}_{h}$ satisfying $a_{\ell}(u-u_{h},v_{h})=0$ for all $v_{h}\in{\cal H}_{h}$,
$\left\|u-u_{h}\right\|_{{\cal H}}\leq C_{2}\Big(1+(kh)^{p}\big\|{\mathcal{A}}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}\Big)\left\|(I-\Pi_{h})u\right\|_{\cal H},$ (4.16)
[…]
$\big\|(I-\Pi_{h}){\mathcal{A}}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}\leq C_{3}\Big(kh+(kh)^{p}\big\|{\mathcal{A}}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}\Big).$ (4.17)
Then, given $0<\varepsilon\leq C_{1}$ there exists $C_{4}>0$ such that if $(kh)^{2p}\big\|{\mathcal{A}}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}\leq\varepsilon$ then
$\inf_{u_{h}\in{\cal H}_{h}\setminus\{0\}}\sup_{v_{h}\in{\cal H}_{h}\setminus\{0\}}\frac{\big|a_{\ell}(u_{h},v_{h})\big|}{\left\|u_{h}\right\|_{{\cal H}}\left\|v_{h}\right\|_{{\cal H}}}\geq\frac{1}{(C_{\rm G1,\ell})^{-1}\Big(1+C_{\rm G2,\ell}(1+C_{2}C_{3}C_{4})\big\|\mathcal{A}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}\Big)}.$ (4.18)
Since the proofs of Corollaries 4.9 and 4.10 are so short, we give them here.
Proof of Corollary 4.9.
By the triangle inequality and the fact that $\|(I-\Pi_{N})u\|_{{\cal H}}\leq\left\|u\right\|_{\cal H}$ (since $\Pi_{N}:{\cal H}\to{\cal H}_{N}$ is the orthogonal projection),
$\left\|u_{N}\right\|_{{\cal H}}\leq\left\|u\right\|_{{\cal H}}+\left\|u-u_{N}\right\|_{{\cal H}}\leq(1+C_{\rm qo,\ell})\left\|u\right\|_{{\cal H}}\leq(1+C_{\rm qo,\ell})\left\|{\mathcal{A}}^{-1}_{\ell}\right\|_{{\cal H}_{0}\to{\cal H}}\left\|f\right\|_{{\cal H}_{0}},$
and the result follows from Lemma 4.8. ∎
Proof of Corollary 4.10.
Given $f\in{\cal H}_{0}$, let $u\in{\cal H}$ be the solution to $a_{\ell}(u,v)=(f,v)_{{\cal H}_{0}}$ for all $v\in{\cal H}$. Now, by the combination of (4.16) and (4.17),
$\left\|u-u_{h}\right\|_{{\cal H}}$ $\leq C_{2}\Big(1+(kh)^{p}\big\|{\mathcal{A}}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}\Big)C_{3}\Big(kh+(kh)^{p}\big\|{\mathcal{A}}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}\Big)\left\|f\right\|_{{\cal H}_{0}}.$
Since $(kh)^{2p}\big\|{\mathcal{A}}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}\leq\varepsilon$ and $\big\|{\mathcal{A}}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}\geq c$, there exists $C_{4}>0$ (depending on $c$ and $\varepsilon$) such that
$\left\|u-u_{h}\right\|_{{\cal H}}\leq C_{2}C_{3}C_{4}\big\|{\mathcal{A}}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}\left\|f\right\|_{{\cal H}_{0}}.$
By the triangle inequality, (4.14) holds with
$C_{\rm sol,\ell}:=(1+C_{2}C_{3}C_{4})\big\|{\mathcal{A}}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}},$
and the result (4.18) follows from Lemma 4.8. ∎
Remark 4.11 (${\mathcal{A}}^{-1}_{\ell}$ vs $({\mathcal{A}}^{*}_{\ell})^{-1}$ in the bound (4.17)).
The bound (4.17) is usually proved with ${\mathcal{A}}^{-1}_{\ell}$ replaced by $({\mathcal{A}}^{*}_{\ell})^{-1}$ (see, e.g., [32, 6, 17]). However, if
[…]
then the definitions of $\mathcal{A}$$^{*}_{\ell}$ and $\mathcal{A}$$_{\ell}$ imply that $({\mathcal{A}}^{*}_{\ell})^{-1}f=\overline{{\mathcal{A}}_{\ell}^{-1}\overline{f}}$, and thus the bound (4.17) is equivalent to the corresponding bound with ${\mathcal{A}}^{-1}_{\ell}$ replaced by $({\mathcal{A}}^{*}_{\ell})^{-1}$. The condition (4.19) holds for all the Helmholtz and Maxwell problems in Lemmas 2.2 and 2.4 either when $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ are Hermitian or when $\mu_{\ell}^{-1}$ and $\epsilon_{\ell}$ corresponding to a radial PML (see [16, Lemma 2.3]).
Furthermore, the proof of [17, Theorem 1.7], which establishes (4.17) with ${\mathcal{A}}^{-1}_{\ell}$ replaced by $({\mathcal{A}}^{*}_{\ell})^{-1}$ for general Helmholtz problems, also proves (4.17).
Remark 4.12 (Link to other results in the literature).
The result that the discrete inf-sup constant is bounded by the continuous inf-sup constant when $\|(I-\Pi_{h})({\mathcal{A}}^{*}_{\ell})^{-1}\|_{{\cal H}_{0}\to{\cal H}}$ is small is [32, Theorem 4.2]. By the Schatz argument [38], [32, Theorem 4.3], if $\|(I-\Pi_{h})({\mathcal{A}}^{*}_{\ell})^{-1}\|_{{\cal H}_{0}\to{\cal H}}$ is small, then the sequence of Galerkin solutions is quasi-optimal, and so [32, Theorem 4.2] is morally equivalent to Corollary 4.9.
Bounding the discrete inf-sup constant for the $h$-FEM applied to each of the Helmholtz and Maxwell problems in §2.
Suppose that the coefficients and domain satisfy the natural regularity requirements for using degree $p$ polynomials; i.e., the domain is $C^{p,1}$ and the coefficients are piecewise $C^{p-1,1}$.
  • All the Helmholtz problems in Lemma 2.2 satisfy the assumptions of Corollary 4.10 by [17, Theorems 1.5 and 1.7] (with [17, Theorem 1.5] giving (4.16) and [17, Theorem 1.7] giving (4.17)).
  • For the Maxwell PEC problem of Part (i) of Lemma 2.4 discretised using Nédélec edge elements of either first or second type, the bound (4.16) under the condition (4.15) is proved in [4, Theorem 1.3]. The bound (4.17) is not proved in [4] (the arguments of [4] use a duality argument that “builds in” the splitting used to prove (4.17)). Therefore, although we expect that the bound (4.14) holds in the preasymptotic regime for the Maxwell PEC problem (via a result analogous to Corollary 4.10), right now [4, Theorem 1.3] only allows us to use Corollary 4.9 in the asymptotic regime (i.e. under the condition that $(kh)^{p}\|{\mathcal{A}}^{-1}_{\ell}\|$ is sufficiently small).
  • As noted above, for the Maxwell impedance problem of Part (ii) of Lemma 2.4, the bound (4.14) is proved directly in [30, Theorem 4.1] for the Maxwell impedance problem in Part (ii) of Lemma 2.4 when $\Omega$ a $C^{2}$ domain, $\mu_{\ell}=\epsilon_{\ell}=I$, $({\cal H}_{N})_{N=0}^{\infty}$ consists of Nédélec edge elements of the second type, and $(kh)^{2}k$ sufficiently small (note that $\|{\mathcal{A}}^{-1}_{\ell}\|\sim k$ here by [26]).
5 Proofs of the main results
Sections 5.1-5.3 are devoted to the proof of Theorem 4.2, and Section 5.4 is devoted to the proof of Lemma 4.8.
5.1 Bounds on $\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\|$ and $\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\|$ in terms of the discrete inf-sup constant
The heart of the proof of Theorem 4.2 is the following result.
Lemma 5.1.
Suppose that the assumptions of §2 hold. Given $N>0$, if exists $0<C_{{\rm dis},N,2}<\infty$ such that
$\inf_{u_{N}\in{\cal H}_{N}\setminus\{0\}}\sup_{v_{N}\in{\cal H}_{N}\setminus\{0\}}\frac{\big|a_{2}(u_{N},v_{N})\big|}{\left\|u_{N}\right\|_{{\cal H}}\left\|v_{H}\right\|_{{\cal H}}}\geq\frac{1}{C_{{\rm dis},N,2}}$ (5.1a)
and
 $\text{ for all }v_{N}\in{\cal H}_{N}\setminus\{0\},\,\sup_{u_{N}\in{\cal H}_{N}\setminus\{0\}}|a_{2}(u_{N},v_{N})|>0,$ (5.1b)
then $\mathsf{A}_{2}^{-1}$ exists and
$\max\Big\{\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{\mathsf{D}},\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{\mathsf{D}^{-1}}\Big\}\leq\Big(\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big)C_{{\rm dis},N,2}.$ (5.2)
Furthermore, if $\mu_{1}=\mu_{2}$, then
$\max\Big\{\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{2},\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{2}\Big\}\leq\frac{m_{+}}{m_{-}}\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}.$ (5.3)
Given Lemma 5.1, to prove Theorem 4.2 we only need to show that the condition (4.4) implies that (5.1) holds with $C_{{\rm dis},N,2}$ bounded above by $\|{\mathcal{A}}^{-1}_{1}\|_{{\cal H}^{*}\to{\cal H}}$.
Proof of Lemma 5.1.
We first show how the bounds on ${\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}$ follow from those on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$. Given a matrix $\mathsf{C}$, let $\mathsf{C}^{\dagger}$ be the conjugate transpose of $\mathsf{C}$ (i.e. the adjoint with respect to $(\cdot,\cdot)_{2}$). Then
$\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{2}=\big\|{\mathsf{I}}-(\mathsf{A}_{2}^{\dagger})^{-1}(\mathsf{A}_{1}^{\dagger})\big\|_{2}.$
Furthermore, by direct calculation,
$\frac{(\mathbf{V}_{1},\mathsf{C}\mathbf{V}_{2})_{\mathsf{D}^{-1}}}{\left\|\mathbf{V}_{1}\right\|_{\mathsf{D}^{-1}}\left\|\mathbf{V}_{2}\right\|_{\mathsf{D}^{-1}}}=\frac{(\mathsf{C}^{\dagger}\mathbf{W}_{1},\mathbf{W}_{2})_{\mathsf{D}}}{\left\|\mathbf{W}_{1}\right\|_{\mathsf{D}}\left\|\mathbf{W}_{2}\right\|_{\mathsf{D}}}\quad\text{ for all }\mathbf{V}_{j}\in\mathbb{C}^{N},$
where $\mathbf{W}_{j}:=\mathsf{D}^{-1}\mathbf{V}_{j}$, $j=1,2,$. Therefore
$\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{\mathsf{D}^{-1}}=\big\|{\mathsf{I}}-(\mathsf{A}_{2}^{\dagger})^{-1}(\mathsf{A}_{1}^{\dagger})\big\|_{\mathsf{D}}.$
Recall from §2 that if $a_{\ell}(\cdot,\cdot)$ satisfies the assumptions in §2, then so does $a^{*}_{\ell}(\cdot,\cdot)$ defined by $a^{*}_{\ell}(u,v)=\overline{a_{\ell}(v,u)}$, with the same constants $C_{\rm G1,\ell}$ and $C_{\rm G2,\ell}$. Furthermore, if the discrete inf-sup condition (4.4) on $a(\cdot,\cdot)$ holds, then it also holds for $a^{*}(\cdot,\cdot)$, with the same $C_{{\rm dis},N,2}$; see, e.g., [25, Lemma 6.5.3], [37, Remark 2.1.45] (this is just the result that the norm of the adjoint operator equals the norm of the original operator).
Therefore, once the bound on $\big\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\big\|_{\mathsf{D}}$ in (4.6) is established, this bound also holds $\mathsf{A}_{2}$ replaced by $\mathsf{A}_{2}^{\dagger}$ and $\mathsf{A}_{1}$ replaced by $\mathsf{A}_{1}^{\dagger}$, so that
$\big\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\big\|_{\mathsf{D}^{-1}}$ $=\big\|{\mathsf{I}}-(\mathsf{A}_{2}^{\dagger})^{-1}(\mathsf{A}_{1}^{\dagger})\big\|_{\mathsf{D}}$
$\leq\Big(\left\|(\mu^{*}_{1})^{-1}-(\mu^{*}_{2})^{-1}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon^{*}_{1}-\epsilon^{*}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big)C_{{\rm dis},N,2}$
$=\Big(\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big)C_{{\rm dis},N,2}.$
Identical reasoning obtains the bound on $\|{\mathsf{I}}-\mathsf{A}_{1}\mathsf{A}_{2}^{-1}\|_{2}$ from that on $\|{\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}\|_{2}$.
We now prove the bounds on ${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}$ in (5.2) and (5.3). By Theorem 3.1, the discrete inf-sup condition (4.4) implies that $\mathsf{A}_{2}:\mathbb{C}^{N}\to\mathbb{C}^{N}$ is invertible. Then, by the definitions of $\mathsf{A}_{\ell}$ (4.1), $\mathsf{S}_{\mu}$, and $\mathsf{M}_{\epsilon}$ (4.2),
${\mathsf{I}}-\mathsf{A}_{2}^{-1}\mathsf{A}_{1}=\mathsf{A}_{2}^{-1}(\mathsf{A}_{2}-\mathsf{A}_{1})=\mathsf{A}_{2}^{-1}\Big(\mathsf{S}_{\mu^{-1}_{2}}-\mathsf{S}_{\mu^{-1}_{1}}-\mathsf{M}_{\epsilon_{2}}+\mathsf{M}_{\epsilon_{1}}\Big)=\mathsf{A}_{2}^{-1}\Big(\mathsf{S}_{\mu^{-1}_{2}-\mu^{-1}_{1}}-\mathsf{M}_{\epsilon_{2}-\epsilon_{1}}\Big).$
Therefore, to prove (5.2), it is sufficient to prove that
$\big\|\mathsf{A}_{2}^{-1}\mathsf{S}_{\mu^{-1}}\big\|_{\mathsf{D}}\leq\left\|\mu^{-1}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}\quad\text{ and }\quad\big\|\mathsf{A}_{2}^{-1}\mathsf{M}_{\epsilon}\big\|_{\mathsf{D}}\leq\left\|\epsilon\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{,}}$
and to prove (5.3), it is sufficient to prove that
$\big\|\mathsf{A}_{2}^{-1}\mathsf{M}_{\epsilon}\big\|_{2}\leq\frac{m_{+}}{m_{-}}\left\|\epsilon\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}.$
It is therefore sufficient to prove that, given $\mathbf{F}\in\mathbb{C}^{N}$, the solutions $\mathbf{U}$ and $\mathbf{W}$ to
$\mathsf{A}_{2}\mathbf{U}=\mathsf{S}_{\mu^{-1}}\mathbf{F}\quad\text{ and }\quad\mathsf{A}_{2}\mathbf{W}=\mathsf{M}_{\epsilon}\mathbf{F}$ (5.4)
satisfy
$\left\|\mathbf{U}\right\|_{\mathsf{D}}\leq\left\|\mu^{-1}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}\left\|\mathbf{F}\right\|_{\mathsf{D}},\quad\left\|\mathbf{W}\right\|_{\mathsf{D}}\leq\left\|\epsilon\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}\left\|\mathbf{F}\right\|_{\mathsf{D}},$ (5.5)
and
$m_{-}\left\|\mathbf{W}\right\|_{2}\leq m_{+}\left\|\epsilon\right\|_{{\cal H}_{0}\to{\cal H}_{0}}C_{{\rm dis},N,2}\left\|\mathbf{F}\right\|_{2}.$ (5.6)
Given $\mathbf{F}\in\mathbb{C}^{N}$, let $\widetilde{f}\in{\cal H}$ and $\widetilde{F}\in{\cal H}^{*}$ be defined by
$\widetilde{f}:=\sum_{j=1}^{N}F_{j}\phi_{j}\quad\text{ and }\quad\widetilde{F}(v)=\big(\mu^{-1}{\mathcal{D}}\widetilde{f},{\mathcal{D}}v\big)_{{\cal H}_{0}}.$ (5.7)
The solutions $u_{N}$ and $w_{N}$ to the variational problems
$a_{2}(u_{N},v_{N})=\widetilde{F}(v_{N})\quad\text{ and }\quad a_{2}(w_{N},v_{N})=(\widetilde{f},v_{N})_{{\cal H}_{0}}\quad\text{ for all }v_{N}\in{\cal H}_{N}$ (5.8)
are then such that
$u_{N}=\sum_{j=1}^{N}U_{j}\phi_{j}\quad\text{ and }\quad w_{N}=\sum_{j=1}^{N}W_{j}\phi_{j},$ (5.9)
with $\mathbf{U}$ and $\mathbf{W}$ the solutions to (5.4).
Now, by the bound (5.1a) involving $C_{{\rm dis},N,2}$, the definition of $u_{N}$ (5.8), the definition of $\widetilde{F}$ (5.7), and the fact that $\left\|{\mathcal{D}}\right\|_{{\cal H}\to{\cal H}_{0}}\leq 1$,
$\left\|u_{N}\right\|_{{\cal H}}\leq C_{{\rm dis},N,2}\sup_{v_{N}\in{\cal H}_{N}}\frac{\big|a_{2}(u_{N},v_{N})\big|}{\left\|v_{N}\right\|_{{\cal H}}}$ $\leq C_{{\rm dis},N,2}\sup_{v_{N}\in{\cal H}_{N}}\frac{\big\|\mu^{-1}{\mathcal{D}}\widetilde{f}\big\|_{{\cal H}_{0}}\left\|{\mathcal{D}}v_{N}\right\|_{{\cal H}_{0}}}{\left\|v_{N}\right\|_{{\cal H}}}$
$\leq C_{{\rm dis},N,2}\left\|\mu^{-1}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\big\|\widetilde{f}\big\|_{{\cal H}}$.
By (1.3), (5.9), and (5.7), $\left\|u_{N}\right\|_{{\cal H}}=\left\|\mathbf{U}\right\|_{\mathsf{D}}$ and $\big\|\widetilde{f}\big\|_{{\cal H}}=\left\|\mathbf{F}\right\|_{\mathsf{D}}$, and the first bound in (5.5) follows. Similarly, by the bound (5.1a) involving $C_{{\rm dis},N,2}$, the definitions of $w_{N}$ (5.8), and the fact that $\left\|\cdot\right\|_{{\cal H}_{0}}\leq\left\|\cdot\right\|_{\cal H}$,
$\left\|w_{N}\right\|_{{\cal H}}\leq C_{{\rm dis},N,2}\sup_{v_{N}\in{\cal H}_{N}}\frac{\big|a_{2}(w_{N},v_{N})\big|}{\left\|v_{N}\right\|_{{\cal H}}}$ $\leq C_{{\rm dis},N,2}\sup_{v_{N}\in{\cal H}_{N}}\frac{\big\|\epsilon\widetilde{f}\big\|_{{\cal H}_{0}}\left\|v\right\|_{{\cal H}_{0}}}{\left\|v\right\|_{{\cal H}}}$
$\leq C_{{\rm dis},N,2}\left\|\epsilon\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\big\|\widetilde{f}\big\|_{{\cal H}_{0}}.$ (5.10)
The second bound in (5.5) then follows since, by (1.3), (5.9), and (5.7), $\left\|w_{N}\right\|_{{\cal H}}=\left\|\mathbf{W}\right\|_{\mathsf{D}}$ and $\|\widetilde{f}\|_{{\cal H}_{0}}\leq\|\widetilde{f}\|_{{\cal H}}=\left\|\mathbf{F}\right\|_{\mathsf{D}}$.
Finally, the bound (5.6) follows from (5.10) by using $m_{-}\|\mathbf{W}\|_{2}\leq\|w_{N}\|_{{\cal H}_{0}}\leq\|w_{N}\|_{{\cal H}}$ and $\|\widetilde{f}\|_{{\cal H}_{0}}\leq m_{+}\|\mathbf{F}\|_{2}$ (with both of these bounds following from (4.3)). ∎
5.2 Norm of solution operator under perturbation
Lemma 5.2 (Norm of solution operator under perturbation).
Suppose that the assumptions of §2 hold. If ${\mathcal{A}}_{1}:{\cal H}\to{\cal H}^{*}$ is invertible and
$\Big(\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big)\|\mathcal{A}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}\leq 1/2,$ (5.11)
then
$\|\mathcal{A}_{2}^{-1}\|_{{\cal H}^{*}\to{\cal H}}\leq 2\|\mathcal{A}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}.$ (5.12)
Proof.
Given $F\in{\cal H}^{*}$, let $u_{2}\in{\cal H}$ be the solution to $a_{2}(u_{2},v)=F(v)$ for all $v\in{\cal H}$. By the definition of $a_{\ell}(\cdot,\cdot)$ (2.1),
$a_{1}(u_{2},v)=F(v)+\big((\mu^{-1}_{1}-\mu^{-1}_{2}){\mathcal{D}}u_{2},{\mathcal{D}}v\big)_{{\cal H}_{0}}-\big((\epsilon_{1}-\epsilon_{2})u_{2},v\big)_{{\cal H}_{0}}.$
By the definition of $\|\mathcal{A}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}$, the fact that $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$, and the bound (5.11),
$\left\|u_{2}\right\|_{{\cal H}}$ $\leq\|\mathcal{A}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}\Big(\left\|F\right\|_{{\cal H}^{*}}+\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\left\|u_{2}\right\|_{{\cal H}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\left\|u_{2}\right\|_{{\cal H}_{0}}\Big)$
$\leq\|\mathcal{A}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}\left\|F\right\|_{{\cal H}^{*}}+\frac{1}{2}\left\|u_{2}\right\|_{{\cal H}},$
and the result (5.12) follows. ∎
Remark 5.3 (Proof of the bound (4.13) under the condition (4.12)).
The condition (4.12) is the same as (5.11), and thus (5.12) holds. Since
[…]
the bound (4.13) follows from (5.12) and the bound
$\left\|{\mathcal{A}}_{2}-{\mathcal{A}}_{1}\right\|_{{\cal H}\to{\cal H}^{*}}\leq\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}};$
this last bound holds since, by the definition of $a_{\ell}(\cdot,\cdot)$ (2.1),
$\big|a_{1}(u,v)-a_{2}(u,v)\big|\leq\Big(\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big)\left\|u\right\|_{\cal H}\left\|v\right\|_{{\cal H}}$
for all $u,v\in{\cal H}$ (since $\|{\mathcal{D}}\|_{{\cal H}\to{\cal H}_{0}}\leq 1$).
Lemma 5.4 (Norm of discrete solution operator under perturbation).
Suppose that the assumptions of §2 hold. Suppose that, with $0<C_{{\rm dis},N,1}<\infty$,
$\inf_{u_{N}\in{\cal H}_{N}\setminus\{0\}}\sup_{v_{N}\in{\cal H}_{N}\setminus\{0\}}\frac{\big|a_{1}(u_{N},v_{N})\big|}{\left\|u_{N}\right\|_{{\cal H}}\left\|v_{H}\right\|_{{\cal H}}}\geq\frac{1}{C_{{\rm dis},N,1}},$ (5.13a)
 $\text{ for all }v_{N}\in{\cal H}_{N}\setminus\{0\},\,\sup_{u_{N}\in{\cal H}_{N}\setminus\{0\}}|a_{1}(u_{N},v_{N})|>0,$ (5.13b)
and
$\Big(\left\|\mu^{-1}_{1}-\mu^{-1}_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}+\left\|\epsilon_{1}-\epsilon_{2}\right\|_{{\cal H}_{0}\to{\cal H}_{0}}\Big)C_{{\rm dis},N,1}\leq 1/2.$ (5.14)
Then (5.1a) and (5.1b) hold with
$C_{{\rm dis},N,2}=2C_{{\rm dis},N,1}.$
Proof.
Let ${\mathcal{A}}_{N,\ell}:{\cal H}_{N}\to({\cal H}_{N})^{*}$, $\ell=1,2,$ be the operators associated to the sesquilinear forms $a_{\ell}(\cdot,\cdot):{\cal H}_{N}\times{\cal H}_{N}\to\mathbb{C}$, $\ell=1,2$. By Theorem 3.1, the condition (5.13) is equivalent to the statement that
$\|{\mathcal{A}}_{N,1}^{-1}\|_{({\cal H}_{N})^{*}\to{\cal H}_{N}}\leq C_{{\rm dis},N,1}.$ (5.15)
The condition (5.14) then implies that (5.11) holds with $\|\mathcal{A}_{1}^{-1}\|_{{\cal H}^{*}\to{\cal H}}$ replaced by $\|{\mathcal{A}}_{N,1}^{-1}\|_{({\cal H}_{N})^{*}\to{\cal H}_{N}}$. The arguments in the proof of Lemma 5.2 then show that
$\big\|{\mathcal{A}}_{N,2}^{-1}\big\|_{({\cal H}_{N})^{*}\to{\cal H}_{N}}\leq 2\big\|{\mathcal{A}}_{N,1}^{-1}\big\|_{({\cal H}_{N})^{*}\to{\cal H}_{N}}.$
Then, by Theorem 3.1 applied to ${\mathcal{A}}_{N,2}$ and (5.15),
$\inf_{u_{N}\in{\cal H}_{N}\setminus\{0\}}\sup_{v_{N}\in{\cal H}_{N}\setminus\{0\}}\frac{\big|a_{2}(u_{N},v_{N})\big|}{\left\|u_{N}\right\|_{{\cal H}}\left\|v_{H}\right\|_{{\cal H}}}\geq\frac{1}{2\big\|{\mathcal{A}}_{N,1}^{-1}\big\|_{({\cal H}_{N})^{*}\to{\cal H}_{N}}}\geq\frac{1}{2C_{{\rm dis},N,1}},$
and the result follows. ∎
5.3 Proof of Theorem 4.2
The assumption (4.4) implies that (5.13) holds with $C_{{\rm dis},N,1}=C_{1}\|{\mathcal{A}}^{-1}_{1}\|_{{\cal H}^{*}\to{\cal H}}$. The condition (4.5) is then (5.14) and thus Lemma 5.4 implies that (5.1a) and (5.1b) hold with
$C_{{\rm dis},N,2}=2C_{1}\|{\mathcal{A}}^{-1}_{1}\|_{{\cal H}^{*}\to{\cal H}}.$
With this value of $C_{{\rm dis},N,2}$, the bounds (5.2) and (5.3) from Lemma 5.1 then become the results (4.6) and (4.7).
5.4 Proofs of Lemma 3.2 and 4.8 (about equivalence of the norms of the continuous and discrete inverses)
We first prove Lemma 3.2, which consists of proving the bounds (3.2) and (3.3).
Proof of (3.2).
We highlight that this argument goes back to at least [2, Lemma 3.4], but since it is short, and crucial to the proof of Lemma 4.8, we include it for completeness. The first inequality in (3.2) is an immediate consequence of the fact that $\|\cdot\|_{{\cal H}^{*}}\leq\|\cdot\|_{{\cal H}_{0}}$ (which follows from the definition of $\|\cdot\|_{{\cal H}^{*}}$ and $\|\cdot\|_{{\cal H}_{0}}\leq\|\cdot\|_{{\cal H}}$). Let $a^{+}_{\ell}(u,v):=a_{\ell}(u,v)+C_{\rm G2,\ell}(u,v)_{{\cal H}_{0}}$ and observe that $a^{+}_{\ell}(\cdot,\cdot)$ is coercive by (2.2). To prove the second inequality in (3.2) it is sufficient to prove that, given $F\in{\cal H}^{*}$, the solution of $a_{\ell}(u,v)=F(v)$ for all $v\in{\cal H}$ satisfies
$\left\|u\right\|_{\cal H}\leq(C_{\rm G1})^{-1}\big(1+C_{\rm G2,\ell}\big\|{\mathcal{A}}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}\big)\left\|F\right\|_{{\cal H}^{*}}.$ (5.16)
Let $u^{\pm}$ by the solutions to
$a^{+}_{\ell}(u^{+},v)=F(v)\quad\text{ and }\quad a_{\ell}(u^{-},v)=C_{\rm G2}(u^{+},v)_{{\cal H}_{0}}\quad\text{ for all }v\in{\cal H};$
these definitions imply that $u=u^{+}+u^{-}$. By (2.2) and the Lax–Milgram lemma,
$\big\|u^{+}\big\|_{{\cal H}}\leq(C_{\rm G1})^{-1}\left\|F\right\|_{{\cal H}^{*}}.$ (5.17)
By the definition of $\|{\mathcal{A}}_{\ell}^{-1}\|_{{\cal H}_{0}\to{\cal H}}$ and the bound (5.17) on $u^{+}$,
$\big\|u^{-}\big\|_{{\cal H}}\leq\|{\mathcal{A}}_{\ell}^{-1}\|_{{\cal H}_{0}\to{\cal H}}C_{\rm G2,\ell}\big\|u^{+}\big\|_{{\cal H}_{0}}\leq\|{\mathcal{A}}_{\ell}^{-1}\|_{{\cal H}_{0}\to{\cal H}}C_{\rm G2,\ell}(C_{\rm G1})^{-1}\left\|F\right\|_{{\cal H}^{*}};$ (5.18)
the bound (5.16) – and hence also the second bound in (3.2) – then follows.
Proof of (3.3).
The first inequality in (3.3) is an immediate consequence of the fact that $\|\cdot\|_{{\cal H}_{0}}\leq\|\cdot\|_{{\cal H}}$. To prove the second inequality, observe that the Gårding-type inequality (2.2) implies that, given $f\in{\cal H}_{0}$,
$C_{\rm G1,\ell}\big\|\mathcal{A}^{-1}_{\ell}f\big\|_{{\cal H}}^{2}$ $\leq C_{\rm G2,\ell}\big\|\mathcal{A}^{-1}_{\ell}f\big\|^{2}_{{\cal H}_{0}}+\big|\langle\mathcal{A}^{-1}_{\ell}f,f\rangle\big|,$
$\leq C_{\rm G2,\ell}\big\|\mathcal{A}^{-1}_{\ell}f\big\|^{2}_{{\cal H}_{0}}+\big\|\mathcal{A}^{-1}_{\ell}f\big\|_{{\cal H}_{0}}\|f\|_{{\cal H}_{0}}.$
Therefore,
$C_{\rm G1,\ell}\big\|\mathcal{A}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}}^{2}\leq C_{\rm G2,\ell}\big\|\mathcal{A}^{-1}_{\ell}\big\|^{2}_{{\cal H}_{0}\to{\cal H}_{0}}+\big\|\mathcal{A}^{-1}_{\ell}\big\|_{{\cal H}_{0}\to{\cal H}_{0}},$
and the second bound in (3.3) follows.
Finally, to prove Lemma 4.8 we observe the argument above proving the second bound in (3.2) remains unchanged when $F\in{\cal H}^{*}$ is replaced by $F\in({\cal H}_{N})^{*}$; indeed, the Lax–Milgram lemma proves that
$\big\|u^{+}_{N}\big\|_{{\cal H}}\leq(C_{\rm G1,\ell})^{-1}\left\|F\right\|_{({\cal H}_{N})^{*}}.$
Furthermore, since $u_{N}^{+}\subset{\cal H}_{N}\subset{\cal H}\subset{\cal H}_{0}$, (4.14) implies that
$\big\|u^{-}_{N}\big\|_{{\cal H}}\leq C_{\rm sol,\ell}C_{\rm G2,\ell}\big\|u^{+}\big\|_{{\cal H}_{0}}\leq C_{\rm sol,\ell}C_{\rm G2,\ell}(C_{\rm G1,\ell})^{-1}\left\|F\right\|_{({\cal H}_{N})^{*}}$
(compare to (5.18)). The result of Lemma 4.8 therefore follows.
Acknowledgements
The author thanks Qiya Hu (Academy of Mathematics and Systems Science, Chinese Academy of Sciences) for asking him the question of whether the results of [18, 22] extend to the time-harmonic Maxwell equations, Théophile Chaumont-Frelet (INRIA, Lille) for discussions about the discrete inf-sup constant of the Helmholtz sesquilinear form, and the referees for many useful comments.